### v2

# Personalizing Agent Privacy Decisions via Logical Entailment

Abstract.
Personal language model-based agents are becoming more widespread for completing tasks on behalf of users; however, this raises serious privacy questions regarding whether these models will appropriately disclose user data. While prior work has evaluated language models on data-sharing scenarios based on general privacy norms, we focus on personalizing language models’ privacy decisions, grounding their judgments directly in prior user privacy decisions. Our findings suggest that general privacy norms are insufficient for effective personalization of privacy decisions. Furthermore, we find that eliciting privacy judgments from the model through In-context Learning (ICL) is unreliable to due misalignment with the user’s prior privacy judgments and opaque reasoning traces, which make it difficult for the user to interpret the reasoning behind the model’s decisions. To address these limitations, we propose ARIEL (Agentic Reasoning with Individualized Entailment Logic), a framework that jointly leverages a language model and rule-based logic for structured data-sharing reasoning. ARIEL is based on formulating personalization of data sharing as an entailment, whether a prior user judgment on a data-sharing request implies the same judgment for an incoming request. Our experimental evaluations on advanced models and publicly-available datasets demonstrate that ARIEL can reduce the F1 score error by 39.1% over language model-based reasoning (ICL), demonstrating that ARIEL is effective at correctly judging requests where the user would approve data sharing. Overall, our findings suggest that combining LLMs with strict logical entailment is a highly effective strategy for enabling personalized privacy judgments for agents.
privacy, agents, large language models, personalization
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
As Personal AI agents increasingly automate tasks, they frequently encounter scenarios requiring the transmission of private user data. Consequently, the agent must determine the appropriateness of such disclosures. Consider the restaurant booking scenario in Figure 1, where a restaurant requests the user’s phone number to finish booking a table. While the agent could consult the user, frequent interruptions lead to ’privacy fatigue,’ causing disengagement and ineffective data control from the user (acquisti2006there; choi2018role). Ideally, to alleviate this cognitive burden, the agent should be capable to autonomously judge when data sharing is appropriate.
Figure 1. High-level overview of a personal agent completing a task on behalf of a user. A third party is requesting user data in order to complete the task. The agent must decide whether it is appropriate to share the user’s information to fulfill the third-party’s request. To assist with the agent’s judgment, the agent is equipped with a knowledge base storing a history of prior user judgments on data-sharing requests.
Prior to Large Language Models (LLMs), Personal Privacy Assistants (PPAs) relied on either rule-based (zhan2022model; zhan2023privacy) or traditional machine learning (ML) approaches (barbosa2019if; amoros2023predicting). These systems leverage a history of user prior user privacy judgments to decide whether to disclose data for incoming requests. For instance (Figure 1), if a user previously authorized sharing a phone number for a restaurant reservation, the agent utilizes this precedent to approve similar future requests. However, these methods face significant limitations: (1) rule-based approaches are domain-specific and difficult to transfer without substantial modification; and (2) ML methods require extensive user data (from many users) to achieve high performance.
The integration of LLMs into agentic workflows (wang2024survey) offers a promising solution to these challenges, as their In-Context Learning (ICL) capabilities enable effective generalization even from sparse data (brown2020language). While recent studies have evaluated privacy judgments by LLMs (mireshghallah2024can; shao2024privacylens), they typically ground their definition of appropriateness in general Contextual Integrity (CI) norms derived from laws or societal consensus (nissenbaum2004privacy; nissenbaum2009privacy). However, relying solely on general norms fails to capture the nuances of user privacy decisions (barkhuus2012mismeasurement; guo2025not; tran2025understanding; zhang2025towards). In the example above, general norms would likely deem sharing a phone number for a restaurant reservation appropriate, yet a user with a restrictive history of data sharing might explicitly reject this request. General norms cannot account for individual privacy decisions, thus being a barrier in the development of personalized privacy agents. This critical gap motivates the central objective of our work.
Objective: Personalizing LLM-based agents to make data-sharing decisions that maximize alignment with prior user decisions, minimize interruptions, and ensure auditability.
To fulfill this objective, we identify three requirements for designing personal LLM agents.
  • Alignment. The agent’s privacy decisions must adhere to the user’s prior privacy judgments, which is important for effective personalization.
  • Interpretability. The agent must provide an interpretable and traceable reasoning for every privacy judgment, allowing the user to audit the process and maintain trust in the agent’s decision-making.
  • User Agency. The agent must preserve user agency by identifying when prior user privacy judgments are insufficient or ambiguous, and escalating such requests to the user rather than forcing a judgment.
We observe that current LLM-based approaches fail to simultaneously satisfy these design requirements. While leveraging ICL on prior user privacy judgments improves personalization over general privacy norms, it does not guarantee alignment with prior user privacy decisions due to limited logical reasoning (lin2025zebralogic) and potential for hallucinations (ji2023survey; kalai2025language; zhang2023siren) from LLMs. Furthermore, the opaque nature of these models undermines interpretability by preventing verification of the reasoning steps used to reach a privacy judgment. Because the LLM-generated reasoning traces are often unfaithful to the actual decision process (barez2025chain; ye2022unreliability), users cannot effectively audit the agent, ultimately compromising their agency.
Thus, we propose ARIEL (Agentic Reasoning with Individualized Entailment Logic), a framework that satisfies the three requirements for personalization (Figure 3). The key insight is framing personalization of privacy judgments as an entailment, where the agent decides whether a user’s judgment on a prior request entails the same judgment for the incoming request. But rather than relying on LLMs to give a final appropriateness judgment, ARIEL combines LLMs reasoning and rule-based components for reliable and aligned reasoning. ARIEL leverages LLMs to generate ontologies that capture semantic relationships (e.g., mapping data sensitivity hierarchies), which are then processed by a rule-based component to provide a final judgment. This framework ensures that the agent acts only when a decision can be inferred by the user’s prior privacy judgments; otherwise, the request is escalated to the user.
We evaluate ARIEL against LLM-based reasoning with general privacy norms and prior user privacy judgments. These evaluations employ state-of-art LLMs – including GPT-5, Gemini 2.5 Pro, and Claude Sonnet 4.5 – on publicly available datasets covering smart home personal assistants (abdi2021privacy) and an educational (shvartzshnaider2016learning) sharing context. We find that (1) grounding privacy judgments in prior user privacy judgments yields significantly better personalization than relying solely on general privacy norms, and (2) ARIEL outperforms pure LLM-based reasoning by reducing the F1 score error of appropriate judgments by up to 39.1%. Our findings indicate that ARIEL is effective at correctly judging requests where the user would approve data sharing. We summarize our contributions below:
  1. (1)
    We demonstrate the limitations of aligning LLM agents using general privacy norms and ICL on prior user privacy judgments to make personalized data-sharing decisions on behalf of users.
  2. (2)
    We propose ARIEL, a framework for grounding the agent’s privacy judgments in prior user privacy judgments by logical entailment with LLM-generated ontologies. Jointly leveraging LLMs and logical entailment overcomes the rigidity of traditional rule-based systems without incurring the inconsistency inherent in pure LLM-based reasoning.
  3. (3)
    We demonstrate that ARIEL outperforms LLM-based reasoning for personalized privacy decisions, while ensuring user agency and interpretability. Notably, we find that performance gains are larger with smaller models, and that ARIEL remains effective with as few as 20 prior requests.
2. Related Work
We review prior work that relate to personalization of LLM-based agents to make privacy decisions aligned with user privacy preferences. We discuss (1) prior work in developing personal privacy assistants and (2) evaluations of privacy norms and personalization in LLMs. Finally, we examine works in (3) neurosymbolic reasoning and their relation to ARIEL.
2.1. Personal Privacy Assistants
Personal Privacy Assistants (PPAs) aim to reduce the user mental burden in making privacy decisions (stover2023investigating). They explore automating mobile app permissions, predicting user privacy preferences, and making information-sharing decisions (morel2025ai). Earlier PPAs targeting mobile app permissions clustered similar user privacy preferences to obtain privacy profiles that help determine mobile app permissions for a user (lin2014modeling; liu2016follow). Follow-up works developed a support vector machine model using a feature set of contextual cues, such as requesting app, the permission type, and the visibility of requesting app (wijesekera2018contextualizing). Other works use vignette surveys that capture contextual factors, such as device type, which serve as features for predicting privacy preferences using traditional ML techniques like logistic regression or collaborative filtering (naeini2017privacy; barbosa2019if; amoros2023predicting). Furthermore, researchers have proposed PPAs that manage data-sharing decisions by decomposing scenarios into Contextual Integrity (CI) information flows. These systems then utilize rule-based (zhan2022model; zhan2023privacy) or ML-based (kokciyan2022taking) approaches to infer the privacy decision. Finally, Natural Language Inference (NLI)-based models have been proposed to detect mismatches between different privacy policies or between policies and user privacy preferences (alshamsan2022detecting; hosseini2021ambiguity; rao2016expecting). While significant progress has been made for PPAs, they remain limited in that they require aggregating all user data for training, which requires users to upload their privacy preferences and prior decisions to a centralized entity. Moreover, these methods struggle with generalizing to different domains.
2.2. Contextual Privacy Awareness in LLMs
Recent work has evaluated LLMs on privacy-decision scenarios derived from general privacy norms, while also evaluating alignment of LLM privacy decisions in personalized settings.
Contextual Integrity (CI)
Recently, researchers studied whether LLM can judge the appropriateness of data-sharing scenarios that are based on CI (nissenbaum2004privacy; nissenbaum2009privacy). In particular, mireshghallah2024can and shao2024privacylens developed ConfAIde and PrivacyLens, respectively, which are datasets that evaluate whether LLMs can reason about sharing data. yi2025privacy studied ambiguity in these datasets to incorporate contextual expansions for improved privacy reasoning. Furthermore, lan2025contextual improved LLM evaluation on PrivacyLens through post-training on a synthetic dataset. More recently, li2025privaci developed a benchmark targeting legal compliance. These datasets encompass privacy norms sourced from either legal statutes or aggregated judgments from crowdsource studies, both of which capture the general public’s privacy expectations on whether certain data sharing is appropriate in the particular context (guo2025not). Other works have operationalized the contextual awareness of LLMs to reduce privacy leakage of user tasks. ghalebikesabi2024operationalizing developed strategies for LLM-based assistants to be CI compliant in form filling tasks. bagdasarian2024airgapagent implemented a minimization module with an LLM to generate a minimized user data set based on the user task. ngong2025protecting utilized LLMs to reduce private information within a user’s prompt. While evaluating LLMs on general privacy norms is important, these norms fail to capture the nuances of individual user preferences (barkhuus2012mismeasurement; guo2025not; tran2025understanding; zhang2025towards). Consequently, reliance on general norms alone is ineffective for developing agents tasked with making personalized privacy decisions.
Personalized Settings
groschupp2025can investigated personalized access control in LLMs, where the LLM must determine whether to grant data access by evaluating an incoming request against the user’s stated access-control preferences (e.g. ”I prefer to not share data unless necessary”). However, it might not be practical to assume that users can explicitly specify access-control preferences. Further, delegating the access control decision to LLMs can struggle to align with the user’s access-control preferences, especially with limitations of LLMs such hallucinations (ji2023survey; kalai2025language; zhang2023siren) and biases (gallegos2024bias). Another work studied how well LLMs align with human preferences in household human-robot interaction contexts (sullivan2025benchmarking). Our work, on the other hand, focuses on a different problem setup, where we evaluate LLMs making privacy decisions on behalf of users. The final work evaluates how well do judgments from LLMs align with user expectations (cheng2024ci). However, their evaluation is quite limited as the dataset is synthetically generated with the author’s annotating the user appropriateness labels.
2.3. Neurosymbolic Approaches for Reasoning
Prior works that are methodologically similar to ours broadly fall under neurosymbolic reasoning, specifically where the LLM acts as an interface to user inputs and offloads the reasoning through rule-based/logical approaches. Such works include using an LLM to translate natural language into a formal intermediate language, which is then offloaded to a deterministic solver to perform the logical reasoning (olausson2023linc; pan2023logic; trinh2024solving; ye2023satlm). While ARIEL does not employ a theorem solver, it does use few-shot prompting with LLMs to map data-sharing requests to levels in a corresponding ontology, which can then use a rule-base component to determine entailment.
3. Personalizing Privacy Decisions
As illustrated in Figure 1, our agent setup involves a personal agent performing tasks on a user’s behalf. The agent has access to the user’s data and manages data-sharing requests from third parties required to complete these tasks. We describe below our problem setup of how the agent decides whether it is appropriate to transmit user data.
3.1. Problem Statement
We address the problem of an LLM-based agent $A_{\theta}$ tasked with making personalized privacy decisions for a user $u$. We assume a non-adversarial setting where the requests $R_{u}$ are not maliciously designed to manipulate the agent $A_{\theta}$. Upon receiving a request $R_{u}$ to share data $d_{u}$, the agent must infer the user’s judgment $l_{u}$ by leveraging a knowledge base $D_{u}$:
$A_{\theta}(D_{u},R_{u}\in\{\text{appropriate, inappropriate}\}.$
The knowledge base $D_{u}$ represents a history of the user’s past privacy judgments $l^{i}_{u}$ on prior requests $R^{i}_{u}$:
(1) $D_{u}=\{P^{i}_{u}\}_{i=1}^{n},\quad\text{where }P^{i}_{u}=(R^{i}_{u},l^{i}_{u}).$
To illustrate the problem setup, consider the following scenario:
Example 3.0.
Let the incoming request $R_{u}$ be a bank asking for the user’s partial Social Security Number (SSN) to open a checking account. The knowledge base $D_{u}$ contains a prior request $R^{i}_{u}$ where the same bank asked for the user’s full SSN for the same purpose, which the user judged as appropriate ($l^{i}_{u}=\texttt{appropriate}$). The agent must decide if it can infer the appropriateness of $R_{u}$ based on this precedent $(R^{i}_{u},l^{i}_{u})$.
We represent each data request as a CI information flow (nissenbaum2004privacy; nissenbaum2009privacy) containing five parameters: (1) data type, the data that is being transmitted; (2) data subject, the owner of the data; (3) data sender, the individual sharing the data; (4) data recipient, the individual receiving the data; (5) transmission principle, the purpose and/or condition under which the data is shared. We formally define a request below:
Definition 3.0 (Request).
Define $R_{u}=\{d_{u},u,s,r,t\}$ as a set containing five parameters where $d_{u}$ is the user’s data, $u$ is the data subject, $s$ is the data sender, $r$ is the data recipient, and $t$ is the transmission principle.
Using Definition 3.2, we can map the prior and incoming request from Example 3.1 to the corresponding five parameters, as the following:
Prior Request data type: full SSN
data subject: user
data sender: agent
data recipient: bank
transmission principle: open checking account
judgment: appropriate
Incoming Request data type: partial SSN
data subject: user
data sender: agent
data recipient: bank
transmission principle: open checking account
judgment:
3.2. Baseline Approaches
Using an LLM to address the privacy decision problem defined above requires addressing a key challenge: how to design the prompt for the LLM. Here, we discuss two baseline approaches to prompt the LLMs to make privacy judgments along with their limitations.
Norm-based Baseline:
The first choice for prompting the LLM is to utilize Privacy Norms. In this approach, the prompt contains three parts. The first part is the instruction to the LLM to make privacy decisions, the second part is the incoming request, and the third part includes the privacy norms. These privacy norms, akin to CI-inspired approaches, can include the general population’s overall acceptability on particular action (fan2024goldcoin). For instance, determining if a user should share a partial SSN with a bank would rely on broad societal acceptability. However, such an approach is limited because general norms lack the granularity required to capture specific, personalized user privacy decisions.
Privacy Preferences-based Baseline:
To address this limitation, an alternative approach is to prompt the LLM explicitly with the user’s privacy preferences. In this approach, the prompt also includes three parts: the general task instruction, the incoming request, and the user’s privacy preferences. These preferences can be based on rules or policies that the user specify, similar to groschupp2025can, or they can include a list of the user’s privacy judgments on prior requests (e.g., leveraging a past decision about sharing a full SSN to infer judgment of incoming request).
Limitations
The prompting approaches discussed above, whether based on general norms or prior user privacy decisions, exhibit limitations in terms of (1) alignment, (2) interpretability, and (3) user agency.
Appropriateness judgments generated directly by LLMs remain unreliable due to limited logical reasoning capabilities (lin2025zebralogic) and the potential for hallucinations that deviate from system prompts (ji2023survey; kalai2025language; zhang2023siren). Consequently, these judgments are not guaranteed to align with user privacy preferences, resulting in erroneous judgments that limit effective personalization.
Furthermore, the decision-making process is opaque. Even when reasoning traces are elicited, they are often unfaithful to the LLM’s actual logic (barez2025chain; ye2022unreliability). This limitation prevents users from accurately interpreting the reasoning behind the agent’s judgment, thereby eroding the user’s trust in the agent.
Finally, because these methods do not align with user preferences and are not inherently interpretable, they hinder user agency. While personalized agents should aim to mitigate privacy fatigue, this objective must not come at the cost of compromising user awareness of and control over how decisions are made on their behalf (mitchell2025fully).
3.3. Requirements
There is a need to develop agents that make privacy decisions while avoiding the limitations mentioned above. In particular, we posit that such agents must satisfy the following requirements:
  1. (1)
    Alignment. The privacy decisions made by the agent on behalf of the users need to accurately reflect the user’s prior privacy decisions. Alignment is crucial for effective personalization of agents.
  2. (2)
    Interpretability. The agent needs to provide a reasoning that is transparent and traceable so that the user stays in the loop throughout the entire privacy-decision process. Interpretability maintains user trust in the system.
  3. (3)
    User Agency. In scenarios where the agent lacks enough information from the user’s prior privacy decisions to make a judgment, the agent must determine when it should escalate the request to the user for assistance. Importantly, the agent needs to accomplish the right tradeoff of automated privacy decision vs user control.
4. ARIEL: Privacy Judgments with Entailment
We propose ARIEL, a system to enable more reliable and structured reasoning for personalized privacy judgments. Below, we describe ARIEL in more detail.
Figure 3. High-level overview of ARIEL, which can be broken down into three components. ① ARIEL generates an ontology for each user based on their prior privacy judgments from the knowledge base using the LLM. ② Once ARIEL receives an incoming request, it goes through each prior request in the user’s knowledge base to determine if entailment holds. It uses the LLM and the generated ontologies to map the parameters in both requests to correspond levels in the ontologies. ③ A set of rules are applied to determine if entailment holds between each mapped prior requests and the mapped prior request. If the incoming request is not entailed by any prior requests, ARIEL escalates the incoming request to the user.
4.1. Overview
4.1.1. Entailment
The core principle of ARIEL is to personalize privacy judgments by determining if a prior judgment made by the user necessitates the same judgment for the incoming request:
$(R^{i}_{u},l^{i}_{u})\implies(R_{u},l^{i}_{u})$ Does the judgment $l^{i}_{u}$ for the prior request $R^{i}_{u}$ logically imply the same judgment for the incoming request $R_{u}$?
This approach addresses the limitations of LLM-based baselines, which do not provide a logical way to infer user judgments from prior decisions. In contrast, entailment treats prior judgments as logical premises that constrain future decisions, offering an interpretable and controllable method to making privacy decisions. This example highlights how entailment enables privacy decision based on prior requests.
Example 4.0.
Referring back to Example 3.1: a user previously approved sharing their full SSN to open a bank account ($l^{i}_{u}=\texttt{appropriate}$). The incoming request asks for a partial SSN in the exact same context. Logically, the user’s approval to share the more sensitive data (full SSN) entails approval to share the less sensitive data (partial SSN).
The main challenge toward applying entailment is the semantic gap between incoming and prior requests, especially if there is no exact lexical overlap between the parameters of the prior and incoming requests. ARIEL addresses this gap using LLM reasoning and rule-based components unified by ontologies. Drawing on prior work in mobile privacy analysis (andow2019policylint; andow2020actions; wang2018guileak), we utilize ontologies to capture entity relationships between parameters in the incoming and prior requests. For example, we can capture the hierarchy of data sensitivity within an ontology where partial SSN is less sensitive than full SSN. However, we simplify the standard graph structure of ontologies into directed paths, where each vertex represents an entity level (detailed in Section 4.2).
4.1.2. Design
The architecture of ARIEL, illustrated in Figure 3, consists of three components:
  1. (1)
    Ontology Generation: use an LLM to generate relevant ontologies for each user over the parameters on incoming and prior requests.
  2. (2)
    Ontology Mapping: map the parameters of the incoming request and relevant prior requests to levels in the generated ontologies.
  3. (3)
    Entailment Logic: apply a set of entailment rules to determine if the user’s prior judgments entail the incoming request.
The three components of ARIEL operate in two phases. In the offline phase, Ontology Generation creates ontologies for each user based on the user’s prior privacy judgments and general privacy norms encoded in the LLM. The online phase processes each incoming request: first, Ontology Mapping selects relevant prior requests with user judgments from the user’s knowledge base, then maps the parameters in each request to the generated ontologies. This step provides a unified representation of the parameters in both the incoming and prior requests, enabling the entailment step. Next, Entailment Logic applies a set of rules to determine if entailment holds between each mapped prior requests and the mapped incoming request. If the incoming request is not entailed by any prior requests, ARIEL escalates the incoming request to the user.
4.1.3. Meeting Requirements
ARIEL satisfies the design requirements outlined in Section 3.3 as follows:
  1. (1)
    Alignment: Appropriateness judgments are grounded strictly in the user’s prior privacy judgments via logical entailment. The prior requests are mapped onto ontologies, which are generated from the user’s prior privacy judgments.
  2. (2)
    Interpretability: All decisions are transparent and auditable. Users can trace ARIEL’s judgment back to the specific prior request and the logical rules used to establish the entailment.
  3. (3)
    User Agency: ARIEL maintains user agency by escalating decisions only when an incoming request is not logically entailed by prior judgments. This prevents the agent from making appropriateness judgments when the user’s prior privacy judgments are insufficient. Furthermore, by providing interpretable and aligned privacy judgments, ARIEL ensures users retain awareness of and control over the automated decision-making process.
4.2. Ontology Generation
We first describe how ARIEL obtains a set of ontologies $O$. The challenge in generating ontologies is that the agent lacks access to all possible values for each parameter in a request, and it is infeasible to anticipate all these values. For example, a data recipient ontology must be sufficiently comprehensive to allow ARIEL to accurately map various recipients within a request to the ontology.
To address this challenge, ARIEL relies on limited initial information: the domain of the incoming request (e.g., an education context) and the user’s prior privacy judgments $D_{u}$. Leveraging these two sources, the LLM generates user-specific ontologies during an offline phase. The LLM’s ability to construct ontologies from limited information enables ARIEL to adapt to new domains without substantial modification.
ARIEL instructs the LLM to generate an ontology for each request parameter, excluding the data subject, which is always the same for each request. The LLM generates the ontology based on a hierarchical relationship. We detail each ontology with its corresponding relationship below (please refer to Appendix A.5 for the exact ontology generation prompts):
  1. (1)
    Data Type. Represents data sensitivity as perceived by the user. Sensitivity increases at higher levels (Level 1 is least sensitive).
  2. (2)
    Data Sender. Represents the sender’s authority to transmit data. Authority decreases at higher levels (Level 1 has most authority).
  3. (3)
    Data Recipient. Represents the user’s trust in the recipient and the recipient’s authority. Trust and authority decrease at higher levels (Level 1 has highest trust/authority).
  4. (4)
    Transmission Principle. Represents the transmission purpose and safeguards. Safeguard strength and purpose relevance decrease at higher levels (Level 1 has strongest safeguards/relevance).
Each hierarchical relationship is derived from two sources: (1) User’s prior privacy judgments: Relationships are inferred directly from the user’s prior privacy judgments ($D_{u}$). For instance, if a user approves a request involving parents but denies an identical request involving friends, ARIEL infers a higher trust level for parents. This ensures the ontology is personalized. (2) General norms: For relationships not present in $D_{u}$, ARIEL leverages the LLM’s encoded societal norms. For example, the LLM can infer that a typical user generally trusts family more than acquaintances. Figure 4 illustrates a data type ontology generated by Gemini 2.5 Pro using this method.
data type Ontology L1. Non-personal, publicly available information.
L2. User preferences and habits that are not directly identifiable.
L3. Information about the user’s home environment and security.
L4. Contact information and social connections.
L5. Sensitive personal and location information.
L6. Highly sensitive health and medical data.
Figure 4. Example of generated data type ontology from Gemini 2.5 Pro.
4.3. Ontology Mapping
In the online phase, ARIEL evaluates an incoming request $R_{u}$ by comparing it against prior requests. The challenge is that the privacy contexts can diverge significantly if requests differ by multiple parameters simultaneously, making direct entailment unreliable. For example, sharing full SSN with a bank to open a checking account is an entirely different context than sharing full SSN with an IRS agent to file taxes. Furthermore, even when requests are similar, the agent requires a structured process for comparing differing parameter values based on some hierarchal relationship in order to determine the entailment (e.g., comparing data recipients of parents vs friends).
To address these challenges, ARIEL employs a structured approach leveraging the generated ontologies $O$, as detailed in Algorithm 1:
Neighboring Requests, To ensure that comparison is done on contextually similar requests, we first select prior requests that ’neighbor’ the incoming request, denoting the neighboring set as $E$. More precisely, $E$ contains prior requests $R^{i}_{u}$ that have a Hamming distance $d$ of at most one with the incoming request $R_{u}$. This is equivalent to ensuring that both requests differ by only one parameter (line 6 in Algorithm 1).
Ontology Mapping. To rigorously compare the differing parameters, we introduce an operator $\leq_{O}$ to formalize the relationship between the parameters of two distinct requests, $R$ and $R^{\prime}$, based on their mapped levels within $O$. Specifically, $R^{\prime}\leq_{O}R$ holds if every parameter in $R^{\prime}$ is ontologically ’less than or equal to’ the corresponding parameter in $R$. We formally define this as:
Definition 4.0 (Ontological Relationship).
Let $R=\{d_{u},u,s,r,t\}$, $R^{\prime}=\{d^{\prime}_{u},u,s^{\prime},r^{\prime},t^{\prime}\}$ be two requests defined in Definition 3.2 that are partially ordered by the set of ontologies $O$. Then $R^{\prime}\leq_{O}R$ is true if $\forall p\in R,\forall p^{\prime}\in R^{\prime}$ either $p^{\prime}<_{o}p$ or $p^{\prime}=_{o}p$ for the relevant ontology $o$.
Then ARIEL iterates through each neighboring prior request $R^{\prime}_{u}$. For each pair, the agent retrieves the differing parameter values, $p$ (from $R$) and $p^{\prime}$ (from $R^{\prime}_{u}$). These values are then mapped to the relevant ontology $o\in O$ (line 12 in Algorithm 1). For this ontological mapping, we use the LLM $\theta$ to determine the relevant ontology $o$ and map $p$ and $p^{\prime}$ to their corresponding levels in $o$
4.4. Entailment Logic
Once the requests are mapped to the ontology, ARIEL applies a set of rules to determine the judgment of $R_{u}$. This involves comparing the mapped levels between the incoming request and each prior request using $\leq_{O}$, and considering the user judgment $l^{i}_{u}$ on the prior request $R^{i}_{u}$. We formally define our rule-based approach for determining entailment below:
Definition 4.0 (Entailment).
Let $R,R^{\prime}$ be two requests with $l^{\prime}$ being an appropriateness label for $R^{\prime}$. And let $O$ be a set of ontologies. We say that the entailment $(R^{\prime},l^{\prime})\implies(R,l^{\prime})$ holds if one of the following two cases hold:
  1. (1)
    $l^{\prime}=\texttt{inappropriate}$ and $R^{\prime}\leq_{O}R$
  2. (2)
    $l^{\prime}=\texttt{appropriate}$ and $R\leq_{O}R^{\prime}$
Algorithm 1 Online phase of ARIEL
1:LLM $\theta$, user $u$, user knowledge base $D_{u}$, incoming request $R_{u}$, set of ontologies $O$
2:appropriate, inappropriate, or undetermined
[…]
5:for $(R^{i}_{u},l^{i}_{u})\in D_{u}$ do
6:  if $d(R^{i}_{u},R_{u})\leq 1$ then $\triangleright$ Requests differ by one parameter
7:   $E=E\bigcup\{(R^{i}_{u},l^{i}_{u})\}$
8:  end if
9:end for
10:for $(R_{u}^{\prime},l^{\prime}_{u})\in E$ do
11:  Let $p\in R_{u}$, $p^{\prime}\in R^{\prime}_{u}$ such that $p\neq p^{\prime}$ $\triangleright$ Differing parameter
12:  $p_{O}\leftarrow\theta(O,p)$, $p^{\prime}_{O}\leftarrow\theta(O,p^{\prime})$ $\triangleright$ Ontological mapping
13:  if $l^{\prime}_{u}==\texttt{inappropriate}\wedge p^{\prime}_{O}\leq p_{O}$ then
14:   $c_{-}\leftarrow c_{-}+1$
15:  else if $l^{\prime}_{u}==\texttt{appropriate}\wedge p_{O}\leq p^{\prime}_{O}$ then
16:   $c_{+}\leftarrow c_{+}+1$
17:  end if
18:end for
19:if $c_{+}>c_{-}$ then $\triangleright$ Determine Judgment
20:  Return appropriate
21:else if $c_{-}>c_{+}$ then
22:  Return inappropriate
23:else
24:  Return Undetermined $\triangleright$ No majority, undetermined
25:end if
ARIEL checks if entailment holds using Definition 4.3 (line 13 in Algorithm 1). We note that multiple neighboring requests can be entailed, with potentially opposing judgments. This could be due to inconsistency in the user’s prior judgments or incomplete information from the ontology. To address this, the agent goes through each neighboring request and keeps a running vote of the appropriate and inappropriate judgments that were entailed. Whichever appropriateness receives the majority vote is the final judgment. If there is no majority, then the agent outputs ’undetermined’ (line 19 in Algorithm 1), signifying that the incoming request should be escalated to the user for further assistance.
Lastly, we note that in scenarios where multiple appropriate and inappropriate prior requests are entailed, ARIEL can be modified for the agent to escalate the incoming request to the user instead of employing majority vote.
5. Evaluation Setup
We present the experimental framework designed to assess the efficacy of ARIEL. In the following sections, we detail the specific datasets used, describe the baselines compared, and define the evaluation metrics employed to quantify performance.
5.1. Datasets for Data-sharing Requests
We use existing datasets that elicit privacy preferences through factorial vignette surveys based on CI. Participants are presented with data-sharing scenarios and asked to indicate their acceptance of each one (abdi2021privacy; apthorpe2018discovering; apthorpe2019evaluating; bhatia2018empirical; hoyle2020privacy; martin2016measuring; shvartzshnaider2016learning). To simulate a user with prior privacy judgments, each of the data-sharing scenarios are converted into requests with appropriateness judgments from the participants. Then, we divided a participant’s set of answered requests into two pairwise disjoint subsets. The first subset was treated as observed prior requests with the user’s answers as the judgments. The second subset represents incoming requests where the participant’s answers were held out as the ground truth for evaluation. Our evaluation was able to employ two such datasets, which we repurposed for our setup as detailed below. See Figure 5 for example requests from both datasets.
5.1.1. Smart Home Personal Assistant (SPA) dataset
The SPA dataset [^1] Dataset can be accessed here: https://osf.io/63wsm/overview (abdi2021privacy) consists of factorial vignette surveys covering different data-sharing scenarios to understand privacy norms in the SPA ecosystem. The scenarios typically involve a voice assistant provider sending user data to individuals/entities who have access to the user’s SPA ecosystem, under a certain purpose and condition. The data subject of a request is always ’user’ and the data sender is always ’assistant provider’. The total number of participants is 1,730, with each participant randomly assigned to answer a subset of the questions. This resulted in approximately 180 different scenarios answered by each participant.
Due to the relatively large number of participant responses from the SPA dataset, we restrict our evaluation to a set of randomly sampled 500 users. To ensure a consistent experimental setup across all users, we standardized the response size by randomly selecting 70 responses per user. For each of the 500 users, we partitioned the responses into 60 prior requests and 10 incoming requests. This yields a total of 5,000 incoming requests, consisting of 1,765 appropriate and 3,235 inappropriate labels. We also perform an ablation study on varying the number of prior requests in Section 6.2.
Lastly, in the dataset, the participant responses are in 1-5 Likert scale where 1-2 correspond to strongly and somewhat unacceptable, 3 corresponds to neutral, and 4-5 correspond to somewhat and strongly acceptable. We convert the responses with 1-2 scores to ”Inappropriate” and the responses with 4-5 to ”Appropriate”. We filter the remaining responses with scores of 3.
5.1.2. Education Dataset.
The Education dataset [^2] Dataset can be accessed here: https://yansh.github.io/papers/HCOMP/ (shvartzshnaider2016learning) consists of factorial vignette surveys involving the transmission of student data within an educational context. For example, a student’s grades are shared by a professor to parents if the student is performing poorly. The data subject of a request is always ’student’. In total, there were 451 respondents, each answering a random subset of the questions which was around 88 and 89.
The participant responses in the dataset contain numerical values, where 1 corresponds to Yes, 2 corresponds to No, and 3-5 correspond to Does not Make Sense options. Specifically, 3 is ’The sender is unlikely to have the information, 2 is ’The receiver would already have the information, and 3 is ’the question is ambiguous.’ We convert responses with 1 to ”Appropriate” and with 2 to ”Inappropriate”. We filter out the remaining responses with 3-5 scores. Lastly, we filter out users that have less than 70 appropriateness judgments, so each user has 60 prior requests and 10 incoming requests. In total, there are 302 users and 30210=3020302*10=3020 total incoming requests, consisting of 1378 appropriate labels and 1642 inappropriate labels.
We present a summary of the statistics of the evaluation datasets after our pre-processing in Table 1.
Dataset # Sampled
Users
# Requests
per User
Total # App # Inapp
SPA 500 10 incoming
60 prior
5000 1765 3235
Education 302 10 incoming
60 prior
3020 1378 1642
Table 1. A summary of the statistics of the evaluation datasets after our pre-processing. This contains the sampled number of users (# Users), sampled number of requests per user (# Requests per User), the total number of incoming requests among all users (Total), the number of incoming requests with appropriate labels (# App), and the number of incoming requests with inappropriate labels (# Inapp).
SPA Dataset data type: email content
data subject: user
data sender: assistant provider
data recipient: partner
transmission principle: reading emails on user’s behalf
Education Dataset data type: grades
data subject: student
data sender: professor
data recipient: parents
transmission principle: if student is performing poorly
Figure 5. Example requests from the SPA and Education Dataset that are provided to the LLMs in our evaluations.
5.2. Baselines
Following the description in Section 3.2, we employ two natural baseline prompt strategies: (1) privacy norms representative of the general population and (2) prior privacy judgments from users.
5.2.1. Norms-based Baseline
Privacy norms capture the general population’s accepted appropriateness of information flows. There are two prompts that we consider.
  1. (1)
    zero-shot. We construct the prompt to utilize the out-of-the-box privacy norm awareness of LLMs, following from recent prior work on evaluating CI-compliance in LLMs (mireshghallah2024can; shao2024privacylens). The prompt directs the LLM to judge the appropriateness of each information flow based on its knowledge of general privacy norms. Prompts can be found in Appendix A.1.
  2. (2)
    privacy norms. We construct the prompt such that the LLM’s privacy judgment is now based on a provided subset of representative privacy norms. These extracted privacy norms are taken directly from the analysis provided in the papers of each dataset. Table 2 contains a subset of the general privacy from both evaluation datasets that were used by the LLMs. The prompt directs the LLM to judge the appropriateness of each information flow based on the provided general privacy norms. Prompts can be found in A.2.
Rule Judgment
SPA Dataset, user recipients
data type: voice command history Inappropriate
data type: bank account details Inappropriate
recipient: neighbors Inappropriate
recipient: visitors in general inappropriate
SPA Dataset, non-user recipients
data type: bank account details Inappropriate
recipient: advertising agency Inappropriate
tp: no purpose/condition Inappropriate
Education Dataset
sender: professor; recipient: graduate schools; data type: grades; tp: with subject’s consent Appropriate
sender: TA; recipient: classmates; data type: grades; tp: subject performing poorly Inappropriate
Table 2. A subset of the top privacy norms for user and non-user recipients from the SPA dataset, and the top appropriate and inappropriate norms from the Education dataset. The LLMs are provided either the general norms from the SPA or education dataset, depending on which dataset is being evaluated. We denote tp as transmission principle.
5.2.2. Privacy Preferences-based Baseline
We provide privacy judgments on prior requests made by a user. There are two prompts that we then consider.
  1. (1)
    In-Context Learning (ICL). We construct the prompt to leverage ICL (brown2020language) by instructing the LLM to infer the user’s privacy preferences from their judgments on prior requests $D_{u}$. When the LLM judges an incoming request $R_{u}$, its decision is therefore personalized based on the inferred preferences in $D_{u}$. These prompts are in Appendix A.3.
  2. (2)
    ICL (w Undet). We modify the instruction in ICL to output ’undetermined’ if the model is unable to confidently judge whether the incoming request is appropriate or inappropriate based on the user’s privacy preferences. These account for scenarios where the user’s privacy preferences may not always provide sufficient information to accurately judge the incoming request. The prompts are in A.4.
Dataset Model zero-shot privacy norms ICL ICL (w Undet) ARIEL
$F1_{\text{A}}$ (%) $F1_{\text{I}}$ (%) $F1_{\text{A}}$ (%) $F1_{\text{I}}$ (%) $F1_{\text{A}}$ (%) $F1_{\text{I}}$ (%) Sup $F1_{\text{A}}$ (%) $F1_{\text{I}}$ (%) Sup $F1_{\text{A}}$ (%) $F1_{\text{I}}$ (%) Sup
SPA Gemini 2.5 Pro 49.4 75.8 59.1 71.5 81.8 91.0 5000 82.4 92.2 4613 87.2 94.2 3797
GPT-5 27.1 78.3 58.4 71.1 79.0 90.6 5000 80.0 91.1 4921 87.4 94.2 3810
Claude Sonnet 4.5 40.0 75.0 57.7 73.6 81.4 91.0 5000 83.3 91.7 4900 86.6 93.8 3824
Gemma 3 4B 45.1 72.8 46.4 73.1 64.1 84.4 5000 62.5 84.2 4999 75.5 88.4 3680
Education Gemini 2.5 Pro 60.1 61.7 38.5 71.6 69.9 80.8 3020 75.5 87.2 2325 84.1 88.2 1681
GPT-5 62.9 55.6 51.2 69.2 74.0 82.9 3020 74.4 83.2 2961 84.8 89.2 1603
Claude Sonnet 4.5 61.4 63.0 62.3 62.5 75.4 82.8 3020 75.0 82.8 2950 84.0 88.4 1608
Gemma 3 4B 64.1 18.0 46.7 66.8 61.7 78.7 3020 63.4 78.8 3020 80.4 86.1 1644
Table 3. Evaluating different methodologies involving LLMs on requests from SPA and Education datasets. We report the F1-score for appropriate ($F1_{\text{A}}$) and inappropriate ($F1_{\text{I}}$) class, and the total number of appropriateness judgments (Sup) for ICL, ICL (w Undet), and ARIEL. We boldface the best $F1_{\text{A}}$ and $F1_{\text{I}}$ scores, respectively, across all methods for each model and dataset. The evaluations show that ARIEL outperforms all other methodologies for all models on both datasets.
Model Not Entailed Entailed
$F1_{\text{A}}$ (%) $F1_{\text{I}}$ (%) Sup $F1_{\text{A}}$ (%) $F1_{\text{I}}$ (%) Sup
SPA Dataset
Gemini 2.5 Pro 69.6 79.4 1203 87.0 94.2 3797
GPT-5 60.4 79.2 1190 85.7 94.0 3810
Claude Sonnet 4.5 68.1 79.8 1176 86.5 94.0 3824
Education
Gemini 2.5 Pro 58.6 74.2 1339 78.7 86.1 1681
GPT-5 65.3 76.4 1417 82.0 88.6 1603
Claude Sonnet 4.5 68.3 76.9 1412 81.8 87.9 1608
Table 4. Analyzing the judgments from ICL on two sets of incoming requests: one where none of the prior requests were entailed by ARIEL (Not entailed), and the other where entailment occurred by ARIEL (Entailed). These results suggest that ARIEL accurately identifies when the appropriateness of incoming requests are difficult to judge based on the user’s prior privacy judgments.
5.3. Models and Evaluation metrics
Models
We evaluate ARIEL and the baselines on three advanced models: Gemini 2.5-Pro (gemini-2.5-pro), GPT-5 (gpt-5-2025-08-07), and Claude Sonnet 4.5 (claude-sonnet-4-5-20250929). Additionally, we include a smaller, open source model, Gemma 3 4B (team2025gemma).
Metrics
We compare the agent’s predicted judgments against the sampled users’ judgments on incoming requests. We report separate F1 scores for both the appropriate and inappropriate classes. Given an evaluation dataset $D$ containing incoming requests $R_{u}$ with judgments $l_{u}$ for each user $u$, we define the precision ($P_{A}$) for the appropriate class ($c^{+}=\texttt{appropriate}$) as:
$P_{A}:=\frac{|(R_{u},l_{u})\in D:A_{\theta}(D_{u},R_{u})=c^{+}\wedge l_{u}=c^{+}|}{|(R_{u},l_{u})\in D:A_{\theta}(D_{u},R_{u})=c^{+}|}.$
We also define recall ($R_{A})$ for the appropriate class as:
$R_{A}:=\frac{|(R_{u},l_{u})\in D:A_{\theta}(D_{u},R_{u})=c^{+}\wedge l_{u}=c^{+}|}{|(R_{u},l_{u})\in D:l_{u}=c^{+}|}.$
Then we can define the F1 score ($F1_{A}$) for the appropriate class as:
$F1_{A}=2\cdot\frac{P_{A}\cdot R_{A}}{P_{A}+R_{A}}.$
A similar process can be done to obtain the F1 Score for the inappropriate class $F1_{I}$.
Evaluation Procedure for general norms. For each unique request in the datasets, we generate a prompt as described in Section 5.2.1. This results in 825 prompts for the SPA dataset and 1,411 for the Education dataset, covering both zero-shot and privacy norms. Once the prompts are generated, we provide all of them to the models to obtain their responses. Then, for each user, we compare their judgment on an incoming request with the LLM’s judgment on that same request, evaluating this on both general norms prompt types.
Evaluation Procedure for user privacy preferences. We generate a prompt for each incoming request with the user’s privacy judgments on prior requests. This results in 5,000 prompts for the SPA dataset and 3020 prompts for the Education dataset, covering ICL, ICL (w Undet), and ARIEL. Incoming requests resulting in an ’undetermined’ judgment are excluded from the evaluation for the respective baselines (i.e., ICL (w Undet) or ARIEL). To account for this partial coverage, we report the Support (Sup) metric, defined as the total number of appropriateness judgments (i.e., ’appropriate’ or ’inappropriate’) made by the model.
6. Results
We present experiments, ablation studies, and error analysis to study the efficacy and robustness of ARIEL. First we evaluate the end-to-end performance improvement of ARIEL over existing baselines in Section 6.1. Then we evaluate how varying the number of prior user judgments affect ARIEL in Section 6.2. Finally, we qualitatively analyze the failure modes of ARIEL in Section 6.3.
6.1. End-to-End Evaluation
We show the results for all approaches in Table 3. Based on these results, we make the following observations:
(1) ARIEL enhances personalized privacy judgments compared to pure LLM reasoning. ARIEL consistently outperforms LLM-based baselines across all datasets and reasoning models, with notable gains in appropriate judgments. For example, ARIEL improves upon ICL (w Undet) by 10.4% $F1_{\text{A}}$ with GPT-5 on the Education dataset. In other words, the $F1_{\text{A}}$ error reduces by 39.1%, demonstrating that logical entailment effectively grounds judgments in prior user privacy judgments. Notably, these improvements are magnified in smaller models. With Gemma 3 4B, ARIEL yields $F1_{\text{A}}$ gains of 13% and 17% over ICL (w Undet) on the SPA and Education datasets, respectively. This demonstrates ARIEL is suitable for on-device deployment of personalized agents.
Figure 6. Ablation study on the number of prior requests with user judgments contained in the user’s knowledge base $D_{u}$. We evaluate ICL w (Undet) and ARIEL on the SPA dataset with Gemini 2.5 Pro. We report the F1-score for appropriate $F1_{\text{A}}$ and inappropriate $F1_{\text{I}}$ class, and the total number of appropriateness judgments (Sup). We find that ARIEL is more robust to varying the number of prior requests compared to ICL w (Undet)
Figure 7. Ablation study on CoT. We evaluate zero-shot, privacy norms, and ICL w (Undet) on the SPA dataset with Gemini 2.5 Pro. We report the F1-score for appropriate $F1_{\text{A}}$ and inappropriate $F1_{\text{I}}$ class. The results suggest that generally CoT has negligible impact on privacy judgments across both models.
(2) ARIEL accurately identifies when the appropriateness of an incoming request is challenging to judge. We observe that the number judgments that were undetermined by ARIEL is higher than ICL (w Undet). For Gemini 2.5 Pro, we see that ARIEL judged 1203 and 1339 incoming requests as undetermined for SPA and Education datasets, respectively, compared to just 387 and 695 for ICL (w Undet). This observation raises the question of whether the appropriateness label of the incoming requests left undetermined by ARIEL can actually be inferred by an LLM based on the user’s prior privacy judgments.
To investigate this question, we analyzed the ICL performance on two subsets of incoming requests: those where ARIEL found logical entailment, and those where it did not (see Table 4). We observe that ICL achieves at least a 10% increase in both $F1_{\text{A}}$ and $F1_{\text{I}}$ on the entailed requests compared to the non-entailed ones. As such, ARIEL accurately identifies requests that are answerable based on prior judgments, effectively escalating only the more challenging ones. Conversely, ICL (w Undet) produces fewer ’undetermined’ outputs, suggesting that LLMs struggle to recognize when they lack sufficient information to make a privacy judgment. This observation aligns with findings in other domains regarding LLMs’ inability to abstain when lacking necessary knowledge (feng2024don; kirichenko2025abstentionbench).
(3) Relying only on the LLM’s awareness of privacy norms is insufficient for personalization. We observe that performance on appropriate judgments from zero-shot is low for both models and datasets. For inappropriate judgments, zero-shot prompting does performs much better than on appropriate judgments, but the results remain suboptimal. Hence, the LLM’s awareness of privacy norms cannot sufficiently capture the nuanced behavior of user privacy preferences within the SPA and Education contexts.
(4) General privacy norms extracted from overall user responses for each dataset can improve performance over zero-shot. We observe that privacy norms improve appropriate judgments for the SPA dataset across all models, and inappropriate judgments for the Education dataset on most. However, including general norms can sometimes degrade performance. For instance, with Gemini 2.5 Pro on the Education dataset, $F1_{\text{A}}$ decreases by $10\%$ while $F1_{\text{I}}$ improves by only $8\%$, suggesting that adding general privacy norms is not always beneficial.
Type Example/Description Total
ICL (w Undet) + CoT
Preference misalignment The user has previously approved sharing contacts with close friends and parents for the purpose of looking up contact details. Since housemates are a similar type of recipient, this request is likely to be considered appropriate 14
Default to inappropriate The user has consistently approved sharing email content with advertising agencies for personalized ads, as long as there are some safeguards in place. Since this request has no safeguards, it is inappropriate. 5
Contradictory Prior Requests The user has previously judged sharing bank account details with business and finance skills for making transactions as inappropriate without any safeguards. Since the current request lacks a transmission principle, it is likely to be considered inappropriate. 1
ARIEL
Ontology Generation An error caused by the hierarchy of the ontology being incorrect. This is due to incomplete information from the user’s privacy preferences. For example, ”housemate” only appears once in the prior request which is not enough knowledge to determine how trustworthy housemate is for the user. 10
Ontology Mapping This typically caused by domain specific issues about the dataset. For SPA, the transmission principle has many combinations for puprose and condition, which can make it hard for the model to map to the correct level. 5
Contradictory Prior Requests The judgments of the prior requests contradict the judgment of the incoming request. 5
Table 5. The different error types of 20 judgments from ICL (w Undet) + CoT and ARIEL on SPA dataset with Gemini 2.5 Pro.
(5) Prior user privacy judgments improve personalized privacy judgments over general privacy norms. By comparing the $F1_{\text{A}}$ and $F1_{\text{I}}$ between ICL and both zero-shot and privacy norms for all datasets and models, it is clear that personalizing LLMs with prior user privacy judgments improves appropriate and inappropriate judgments over general privacy norms. The improvement over general privacy norms can in some cases be as large as 20% for $F1_{\text{A}}$.
(6) Personalized privacy judgments improve when including undetermined as an option. When comparing ICL and ICL (w Undet), we observe an improvement in $F1_{\text{A}}$ and $F1_{\text{I}}$ for all state-of-art reasoning models. These results suggest that a hybrid approach of combining automated agent judgments with occasional escalation to the user can be highly effective for personalization.
6.2. Ablation Evaluation
We evaluate how the number of prior request affect the performance of ARIEL and baseline approaches. Moreover, we investigate how the choice of reasoning affects LLM privacy judgments.
6.2.1. Number of prior requests
We explore how varying the number of prior requests with user judgments affect the agent’s judgments. Specifically, we compare the $F1_{A}$, $F1_{I}$, and Support (Sup) of ARIEL against ICL (w Undet) on the SPA dataset using Gemini 2.5 Pro. As evident in Figure 6, while the $F1_{I}$ gap remains relatively constant, the disparity in $F1_{A}$ widens significantly as the number of prior requests decreases. Notably, when reducing the prior requests from 60 to 20, $F1_{A}$ drops by only 3.4% for ARIEL, compared to an 8.8% decline for ICL (w Undet). This demonstrates that ARIEL is more robust to limited prior privacy judgments. Nonetheless, we observe that the number of entailments generated by ARIEL decreases substantially as the number of prior requests decreases. This result is expected, as there are fewer candidate requests available to satisfy the entailment criteria (i.e., fewer prior requests differ by exactly one parameter from the incoming request).
6.2.2. Assessing impact of reasoning with Chain of Thought (CoT)
To assess the impact of reasoning strategy choice on appropriateness judgments, we implement CoT by prompting the LLM to provide a reasoning before making a judgment (wei2022chain). Inspired by findings that CoT affects privacy leakage (mireshghallah2024can), we evaluate its effect on judgment accuracy across zero-shot, privacy norms, and ICL (w Undet). Figure 7 indicates that CoT typically provides negligible changes in performance. However, we observe for zero-shot with Gemini 2.5 Pro that the $F1_{A}(\%)$ and $F1_{I}(\%)$ is significantly impacted by CoT, both positively and negatively.
While CoT provides mixed results for general privacy norms, the results clearly demonstrate that adding CoT to both ICL and ICL (w Undet) consistently hurts their performance. When manually inspecting the generated reasoning traces, we notice that the LLM makes misaligned assertions about the user’s privacy judgments, which propagate all the way to the model’s final judgment (see Section 6.3 for more information). This result also demonstrates that eliciting explainability from LLM-based reasoning can degrade the appropriateness judgment. In contrast, ARIEL provides a clear, traceable reasoning while maintaining well-grounded personalized privacy judgments.
6.3. Qualitative Analysis
Finally, we qualitatively analyze incorrect judgments made by both ICL (w Undet) + CoT and ARIEL. We sample 20 examples from each method on the SPA dataset with Gemini 2.5 Pro. The results are detailed in Table 5.
6.3.1. Error Analysis on incorrect appropriateness judgments for ICL (w Undet) + CoT
Analysis of the reasoning traces reveals that Gemini 2.5 Pro occasionally performs implicit entailment when identifying prior requests that share identical parameters except for one. This behavior is noteworthy, as we did not explicitly instruct the model to use this logic, but only to extract patterns from user privacy judgments. We find that the majority of errors stem from preference misalignment (14 examples), where the model draws incorrect equivalencies between distinct parameter values to do this entailment. For instance, given prior requests involving ’close friends’ or ’parents’ as recipients and an incoming request involving ’housemates,’ the model infers that the privacy judgment should remain consistent due to perceived similarity in type of recipient. However, the user may not view these three recipients as similar.
We also attribute a subset of errors to the model’s inherent conservatism regarding information sharing, which we describe as default to inappropriate (5 examples). This occurs when the model concludes it cannot use the most similar prior requests that are judged as ’appropriate’ by the user. But rather than outputting ’undetermined’, the model reasons that the judgment is inappropriate. Finally, we identified one instance of contradictory prior requests, where the provided user history contained judgments that logically conflicted with the ground truth for the incoming request.
6.3.2. Error analysis on incorrect appropriateness judgments for ARIEL
We classify errors in ARIEL into three categories: (1) ontology generation, (2) ontology mapping, and (3) contradictory prior requests. The majority of errors stem from ontology generation, where the generated ontology produces incorrect entailments (10 examples). This typically arises from data sparsity when specific parameter values (e.g., ’housemates’ or ’close friends’) appear only once or twice in the user’s prior requests. The model struggles to accurately reflect the user’s privacy preferences onto the ontology. A second source of error is ontology mapping, where the model fails to map parameter values to the ontology correctly (5 examples). This is often caused by domain-specific idiosyncrasies in the evaluation dataset. For instance, the transmission principle has specific format combining both purpose and condition, which can be challenging for the model to map. Finally, the remaining errors (5 examples) result from contradictory prior requests. For instance, there was a prior request where user had judged sharing bank account details with business and finance skills for the purpose of making transactions as inappropriate. The incoming request is identical as the prior, except there is no purpose. Thus, the incoming request should also be inappropriate, since removing the purpose likely makes the request more inappropriate. Both ICL (w Undet) +CoT and ARIEL judged this request inappropriate, however, the user judged the incoming request to be appropriate.
7. Discussion
Based on the results and takeaways, we discuss in more details the broad implications of this work as well as its limitations.
7.1. Personalization and Privacy Norms
Our results in Section 6 suggest that effective personalization of LLM agents requires grounding judgments in prior user privacy judgments, as general norms are insufficient. In this paper, we consider user judgments to prior requests as proxies for the user’s privacy preferences. However, this finding does not invalidate use general norms in LLMs. Aligning models with relevant contextual norms remains crucial for many data-sharing scenarios (mireshghallah2024can; shao2024privacylens). Our work highlights that when personalization is a crucial component for making privacy decisions, then the LLM’s privacy judgments should be grounded in user-specific preferences rather than general privacy norms.
7.2. Personalization Challenges
Personalizing LLM-based agents introduces several challenges that must be considered.
Contradictory privacy preferences. The first challenge, highlighted by the error analysis in Section 6.3, arises from contradictions within the user’s own history of prior judgments. For example, the user says it is inappropriate to share their banking information with a third party if their data was confidential, but deems it appropriate if there is no safeguard on the data. In such scenarios, the agent may execute the entailment logic correctly but still reach an incorrect privacy judgment. Consequently, the agent must (1) detect these user privacy judgment contradictions and (2) resolve them, typically by escalating the request to the user. The fundamental difficulty lies in distinguishing whether an incorrect privacy judgment stems from an error in the entailment process or an inconsistency in the user’s prior privacy judgments.
Evolving privacy preferences. The second challenge involves the dynamic nature of privacy preferences as they can change overtime. For example, a user can become more conservative about sharing their data as they get older. Hence, a system similar to ARIEL requires a mechanism to detect when the user’s privacy preferences are shifting and determine which preferences must be updated. One potential mechanism could trade-off exploration and exploitation by escalating incoming requests to the user every once in a while to confirm the agent’s decision.
7.3. User Agency in Personal AI Agents
While personalized agents reduce the burden of privacy management, they must not remove user agency (mitchell2025fully). This observation presents a tradeoff where the agent automates clear decisions but escalates cases where the judgment cannot be easily inferred by the user’s prior judgments. ARIEL implements this via an ’undetermined’ option, and our results confirm that identifying these non-inferable cases drives ARIEL’s judgment accuracy gains. Furthermore, by providing transparent reasoning traces for user auditing, ARIEL successfully balances automation with necessary human oversight. Future work could operationalize this transparency through a dedicated user interface, enabling users to visualize the agent’s logic and intervene in decision-making when necessary.
7.4. Limitations
There are three main limitations involving our evaluation in this work. The first is that our evaluation datasets simulate prior requests by partitioning the users requests into two pairwise disjoint sets, where one set stores already observed requests while the other contains new requests. While we believe this reasonably approximates our problem setup, it does not incorporate user decision-making over time. Unfortunately, to our knowledge, there are no existing publicly-available datasets that incorporates user privacy preferences over time. An interesting future work could elicit user privacy preferences over a period of time, rather than giving a one-time questionnaire to participants. The second is our evaluation does not consider cold-start cases where the user has not made prior judgments on requests– essentially, no privacy preferences. While falling back to general privacy norms is a potential solution, our results indicate they are often insufficient for personalized privacy decisions. An alternative solution could be to develop minimal question sets to elicit the user’s privacy preferences (guo2025not; tran2025understanding; zhang2025towards). However, as our work focuses on inferring privacy judgments on behalf of users, we leave this elicitation as a future work. Lastly, our work assumes a non-adversarial setting where the incoming requests are not maliciously designed. While this limits applicability in open environments, it is appropriate for controlled ecosystems where input sources are authenticated and vetted.
8. Conclusion
In this work, we explored personalization of LLM-based agents for making privacy judgments. We identified that general privacy norms are insufficient for effective personalization, and LLM-based reasoning with ICL does not guarantee judgments faithfully grounded on user’s prior privacy judgments. We addressed these limitations by designing ARIEL to ensure judgments are grounded in user’s prior judgments with clear and traceable reasoning that the user can reference. Although ARIEL is intended to accurately make privacy judgments on behalf of users, it is designed with user agency as a key component of the framework. Our evaluations demonstrate that ARIEL can improve appropriate judgments by reducing the F1 score error of appropriate judgments by up to 39.1% over LLM reasoning. Overall, our findings suggest that jointly leveraging LLMs with logical entailment is a promising direction for designing agentic frameworks for personalized privacy judgments.
Acknowledgements.
We would like to thank Adrià Gascón, Sarah Meiklejohn, Eugene Bagdasarian, Daniel Ramage, Zhichen Zhao, and Ku Kogan for their constructive discussions and support.
Appendix A Prompts for Evaluation
We provide the set of prompts that were used in our experimental evaluations below.
A.1. Zero-shot Prompts
Prompt 1: Zero-shot prompt for SPA dataset Prompt 2: Zero-shot prompt for Education dataset
A.2. Privacy Norms Prompts
Prompt 3: Privacy norms prompt for SPA dataset Prompt 4: Privacy norms prompt for Education Dataset
A.3. ICL Prompts
Prompt 5: ICL prompt SPA dataset Prompt 6: ICL prompt for Education dataset
A.4. ICL (w Undet) Propmts
Prompt 7: ICL (w Undetermined) prompt for SPA dataset Prompt 8: ICL (w Undetermined) prompt for Education dataset
A.5. ARIEL Prompts
Prompt 9: Ontology generation prompt for SPA dataset Prompt 10: Ontology generation prompt for Education dataset Prompt 11: Ontology mapping prompt

### v3

@@
 Abstract.
-Personal language model-based agents are becoming more widespread for completing tasks on behalf of users; however, this raises serious privacy questions regarding whether these models will appropriately disclose user data.
-While prior work has evaluated language models on data-sharing scenarios based on general privacy norms, we focus on personalizing language models’ privacy decisions, grounding their judgments directly in prior user privacy decisions.
-Our findings suggest that general privacy norms are insufficient for effective personalization of privacy decisions.
-Furthermore, we find that eliciting privacy judgments from the model through In-context Learning (ICL) is unreliable to due misalignment with the user’s prior privacy judgments and opaque reasoning traces, which make it difficult for the user to interpret the reasoning behind the model’s decisions.
-To address these limitations, we propose ARIEL (Agentic Reasoning with Individualized Entailment Logic), a framework that jointly leverages a language model and rule-based logic for structured data-sharing reasoning.
-ARIEL is based on formulating personalization of data sharing as an entailment, whether a prior user judgment on a data-sharing request implies the same judgment for an incoming request.
-Our experimental evaluations on advanced models and publicly-available datasets demonstrate that ARIEL can reduce the F1 score error by 39.1% over language model-based reasoning (ICL), demonstrating that ARIEL is effective at correctly judging requests where the user would approve data sharing.
-Overall, our findings suggest that combining LLMs with strict logical entailment is a highly effective strategy for enabling personalized privacy judgments for agents.
+Personal large language model (LLM) agents increasingly perform tasks that require access to user data, raising concerns about appropriate data disclosure.
+We show that relying solely on LLMs to make data-sharing decisions is insufficient.
+Prompting LLMs with general privacy norms fails to capture individual users’ privacy preferences, while providing prior user data-sharing decisions through in-context learning (ICL) leads to unreliable and opaque reasoning.
+To address these limitations, we propose ARIEL (Agentic Reasoning with Individualized Entailment Logic), a framework that combines LLMs with rule-based logic to enable structured, personalized privacy reasoning.
+The core mechanism of ARIEL determines whether a user’s prior decision on a data-sharing request logically entails the same decision for a new request.
+Experimental evaluations using advanced models and public datasets show that ARIEL reduces the F1 error rate for appropriate judgments by 40.6% compared to standard ICL-based reasoning, indicating that ARIEL is effective at correctly judging requests where the user would approve data sharing.
+These results demonstrate that integrating LLMs with logical entailment provides an effective and interpretable approach for automating personalized privacy decisions.
 privacy, agents, large language models, personalization
@@
 Consequently, the agent must determine the appropriateness of such disclosures.
-Consider the restaurant booking scenario in Figure 1, where a restaurant requests the user’s phone number to finish booking a table.
-While the agent could consult the user, frequent interruptions lead to ’privacy fatigue,’ causing disengagement and ineffective data control from the user (acquisti2006there; choi2018role).
-Ideally, to alleviate this cognitive burden, the agent should be capable to autonomously judge when data sharing is appropriate.
+Consider the restaurant booking scenario in Figure 1, where a restaurant requests the user’s phone number to complete the table reservation.
+While the agent could seek the user’s permission to transmit their phone number, frequent interruptions lead to ’privacy fatigue,’ causing disengagement and ineffective data control by the user (acquisti2006there; choi2018role).
+To alleviate this cognitive burden, the agent should ideally be capable of autonomously judging when data sharing is appropriate.
+Prior to Large Language Models (LLMs), Personal Privacy Assistants (PPAs) relied on either rule-based (zhan2022model; zhan2023privacy) or traditional machine learning (ML) approaches (barbosa2019if; amoros2023predicting).
+These systems leverage a history of prior user privacy judgments to decide whether to disclose data for incoming requests.
+For instance, as shown in Figure 1, if a user previously authorized sharing a phone number for a restaurant reservation, the agent can utilize this precedent to approve similar future requests.
+However, these methods face significant limitations: (1) rule-based approaches are domain-specific and difficult to transfer without substantial modification; and (2) ML methods require extensive user data (from many users) to achieve high performance.
 Figure 1.
-High-level overview of a personal agent completing a task on behalf of a user.
-A third party is requesting user data in order to complete the task.
-The agent must decide whether it is appropriate to share the user’s information to fulfill the third-party’s request.
-To assist with the agent’s judgment, the agent is equipped with a knowledge base storing a history of prior user judgments on data-sharing requests.
-Prior to Large Language Models (LLMs), Personal Privacy Assistants (PPAs) relied on either rule-based (zhan2022model; zhan2023privacy) or traditional machine learning (ML) approaches (barbosa2019if; amoros2023predicting).
-These systems leverage a history of user prior user privacy judgments to decide whether to disclose data for incoming requests.
-For instance (Figure 1), if a user previously authorized sharing a phone number for a restaurant reservation, the agent utilizes this precedent to approve similar future requests.
-However, these methods face significant limitations: (1) rule-based approaches are domain-specific and difficult to transfer without substantial modification; and (2) ML methods require extensive user data (from many users) to achieve high performance.
+A schematic visualization of the problem setup with ARIEL, our proposed privacy-reasoning framework, operating in a larger end-to-end agentic system.
+① A user sends a task to the agent to complete.
+② A third party requests user data in order to complete the task.
+③ After the data request has been authenticated, the agent retrieves the relevant user data and prior user judgments on data-sharing requests from memory.
+④ The agent decides whether it is appropriate to share the user’s information to fulfill the third party’s request.
+ARIEL either allows or denies transmitting the user data, or escalates the data-sharing request to the user for further assistance.
 The integration of LLMs into agentic workflows (wang2024survey) offers a promising solution to these challenges, as their In-Context Learning (ICL) capabilities enable effective generalization even from sparse data (brown2020language).
@@
 However, relying solely on general norms fails to capture the nuances of user privacy decisions (barkhuus2012mismeasurement; guo2025not; tran2025understanding; zhang2025towards).
-In the example above, general norms would likely deem sharing a phone number for a restaurant reservation appropriate, yet a user with a restrictive history of data sharing might explicitly reject this request.
+In the example above, it is generally acceptable to share a phone number for a restaurant reservation, yet a user with a restrictive history of data sharing might explicitly reject this request.
 General norms cannot account for individual privacy decisions, thus being a barrier in the development of personalized privacy agents.
 This critical gap motivates the central objective of our work.
-Objective: Personalizing LLM-based agents to make data-sharing decisions that maximize alignment with prior user decisions, minimize interruptions, and ensure auditability.
+Objective: Personalizing LLM-based agents to make data-sharing decisions that maximize alignment with prior user decisions, minimize user intervention, and ensure auditability.
 To fulfill this objective, we identify three requirements for designing personal LLM agents.
   • Alignment. The agent’s privacy decisions must adhere to the user’s prior privacy judgments, which is important for effective personalization.
-  • Interpretability.
-The agent must provide an interpretable and traceable reasoning for every privacy judgment, allowing the user to audit the process and maintain trust in the agent’s decision-making.
+  • Interpretability. The agent must provide interpretable and traceable reasoning for every privacy judgment, allowing the user to audit the process and maintain trust in the agent’s decision-making.
   • User Agency.
-The agent must preserve user agency by identifying when prior user privacy judgments are insufficient or ambiguous, and escalating such requests to the user rather than forcing a judgment.
+The agent must preserve user agency by identifying when prior user privacy judgments are insufficient or ambiguous then escalating such requests to the user rather than forcing a judgment.
 We observe that current LLM-based approaches fail to simultaneously satisfy these design requirements.
-While leveraging ICL on prior user privacy judgments improves personalization over general privacy norms, it does not guarantee alignment with prior user privacy decisions due to limited logical reasoning (lin2025zebralogic) and potential for hallucinations (ji2023survey; kalai2025language; zhang2023siren) from LLMs.
+While leveraging ICL on prior user privacy judgments improves personalization over general privacy norms, it does not achieve effective alignment with prior user privacy decisions due to limited logical reasoning (lin2025zebralogic) and potential for LLM hallucinations (ji2023survey; kalai2025language; zhang2023siren).
 Furthermore, the opaque nature of these models undermines interpretability by preventing verification of the reasoning steps used to reach a privacy judgment.
-Because the LLM-generated reasoning traces are often unfaithful to the actual decision process (barez2025chain; ye2022unreliability), users cannot effectively audit the agent, ultimately compromising their agency.
-Thus, we propose ARIEL (Agentic Reasoning with Individualized Entailment Logic), a framework that satisfies the three requirements for personalization (Figure 3).
-The key insight is framing personalization of privacy judgments as an entailment, where the agent decides whether a user’s judgment on a prior request entails the same judgment for the incoming request.
-But rather than relying on LLMs to give a final appropriateness judgment, ARIEL combines LLMs reasoning and rule-based components for reliable and aligned reasoning.
-ARIEL leverages LLMs to generate ontologies that capture semantic relationships (e.g., mapping data sensitivity hierarchies), which are then processed by a rule-based component to provide a final judgment.
-This framework ensures that the agent acts only when a decision can be inferred by the user’s prior privacy judgments; otherwise, the request is escalated to the user.
-We evaluate ARIEL against LLM-based reasoning with general privacy norms and prior user privacy judgments.
-These evaluations employ state-of-art LLMs – including GPT-5, Gemini 2.5 Pro, and Claude Sonnet 4.5 – on publicly available datasets covering smart home personal assistants (abdi2021privacy) and an educational (shvartzshnaider2016learning) sharing context.
-We find that (1) grounding privacy judgments in prior user privacy judgments yields significantly better personalization than relying solely on general privacy norms, and (2) ARIEL outperforms pure LLM-based reasoning by reducing the F1 score error of appropriate judgments by up to 39.1%.
+Because LLM-generated reasoning traces are often unfaithful to the actual decision process (barez2025chain; ye2022unreliability), users cannot effectively audit the agent, ultimately compromising their agency.
+Thus, we propose ARIEL (Agentic Reasoning with Individualized Entailment Logic), a framework that satisfies these three requirements.
+Our key insight is to frame personalized privacy judgments as logical entailment: determining whether a user’s judgment on a prior request entails the same judgment for an incoming request.
+Rather than relying on LLMs to provide an opaque appropriateness judgment, ARIEL combines LLM-based reasoning and rule-based components for reliable and aligned judgments.
+Specifically, ARIEL leverages an LLM to generate ontologies that capture semantic relationships from the user’s prior privacy judgments (e.g., a data sensitivity hierarchy), which are then processed by a rule-based component to provide a final judgment.
+This framework ensures that the agent acts only when a decision can be inferred from the user’s prior privacy judgments; otherwise, the request is escalated to the user (Figure 1).
+We evaluate ARIEL against LLM-based reasoning baselines that rely on either general privacy norms or prior user privacy judgments.
+These evaluations were conducted using state-of-the-art LLMs—including GPT-5, Gemini 2.5 Pro, and Claude Sonnet 4.5—on publicly available datasets covering smart home assistant (abdi2021privacy) and educational (shvartzshnaider2016learning) data-sharing contexts.
+We find that (1) grounding privacy judgments in prior user privacy judgments yields significantly better personalization than relying solely on general privacy norms, and (2) ARIEL outperforms pure LLM-based reasoning by reducing the F1 score error of appropriate judgments by up to 40.6%.
 Our findings indicate that ARIEL is effective at correctly judging requests where the user would approve data sharing.
@@
   1. (1)
-    We demonstrate the limitations of aligning LLM agents using general privacy norms and ICL on prior user privacy judgments to make personalized data-sharing decisions on behalf of users.
+    We demonstrate the limitations of aligning LLM agents with general privacy norms, as well as the shortcomings of current LLM-based approaches that use prior user judgments for personalized data-sharing decisions.
   2. (2)
-    We propose ARIEL, a framework for grounding the agent’s privacy judgments in prior user privacy judgments by logical entailment with LLM-generated ontologies.
+    We propose ARIEL, a framework for grounding the agent’s privacy judgments in prior user privacy judgments through logical entailment with LLM-generated ontologies.
 Jointly leveraging LLMs and logical entailment overcomes the rigidity of traditional rule-based systems without incurring the inconsistency inherent in pure LLM-based reasoning.
   3. (3)
-    We demonstrate that ARIEL outperforms LLM-based reasoning for personalized privacy decisions, while ensuring user agency and interpretability.
-Notably, we find that performance gains are larger with smaller models, and that ARIEL remains effective with as few as 20 prior requests.
+    We show that ARIEL outperforms LLM-based reasoning for personalized privacy decisions, while incorporating user agency and interpretability.
+Notably, we find that performance gains are larger with smaller models (Gemma 3 4B) and that ARIEL remains effective with as few as 20 prior requests.
 2. Related Work
-We review prior work that relate to personalization of LLM-based agents to make privacy decisions aligned with user privacy preferences.
-We discuss (1) prior work in developing personal privacy assistants and (2) evaluations of privacy norms and personalization in LLMs.
+We review closely-related work on personalizing LLM-based agents to make privacy decisions aligned with user privacy preferences.
+We discuss (1) prior work in developing personal privacy assistants and (2) evaluations of privacy norms and user preferences in LLMs.
 Finally, we examine works in (3) neurosymbolic reasoning and their relation to ARIEL.
 2.1. Personal Privacy Assistants
-Personal Privacy Assistants (PPAs) aim to reduce the user mental burden in making privacy decisions (stover2023investigating).
+Personal Privacy Assistants (PPAs) aim to reduce the user mental burden in making privacy decisions (colnago2020informing; stover2023investigating).
 They explore automating mobile app permissions, predicting user privacy preferences, and making information-sharing decisions (morel2025ai).
-Earlier PPAs targeting mobile app permissions clustered similar user privacy preferences to obtain privacy profiles that help determine mobile app permissions for a user (lin2014modeling; liu2016follow).
-Follow-up works developed a support vector machine model using a feature set of contextual cues, such as requesting app, the permission type, and the visibility of requesting app (wijesekera2018contextualizing).
-Other works use vignette surveys that capture contextual factors, such as device type, which serve as features for predicting privacy preferences using traditional ML techniques like logistic regression or collaborative filtering (naeini2017privacy; barbosa2019if; amoros2023predicting).
-Furthermore, researchers have proposed PPAs that manage data-sharing decisions by decomposing scenarios into Contextual Integrity (CI) information flows.
+Earlier PPAs targeting mobile app permissions clustered similar user privacy preferences to derive privacy profiles that help determine mobile app permissions for individual users (lin2014modeling; liu2016follow).
+Follow-up works utilized contextual factors, such as device type, requesting app, and the permission type, as features for predicting privacy preferences using traditional ML techniques like logistic regression or collaborative filtering (naeini2017privacy; barbosa2019if; amoros2023predicting; wijesekera2017feasibility; wijesekera2018contextualizing).
+Other works have proposed PPAs that manage data-sharing decisions by decomposing scenarios into Contextual Integrity (CI) information flows.
 These systems then utilize rule-based (zhan2022model; zhan2023privacy) or ML-based (kokciyan2022taking) approaches to infer the privacy decision.
@@
 2.2. Contextual Privacy Awareness in LLMs
-Recent work has evaluated LLMs on privacy-decision scenarios derived from general privacy norms, while also evaluating alignment of LLM privacy decisions in personalized settings.
-Contextual Integrity (CI)
-Recently, researchers studied whether LLM can judge the appropriateness of data-sharing scenarios that are based on CI (nissenbaum2004privacy; nissenbaum2009privacy).
+Recent work has evaluated LLMs on privacy-decision scenarios derived from general privacy norms, while also evaluating alignment of LLM privacy decisions with user privacy preferences.
+General Privacy Norms
+Recently, researchers studied whether LLMs can judge the appropriateness of data-sharing scenarios that are based on CI (nissenbaum2004privacy; nissenbaum2009privacy).
 In particular, mireshghallah2024can and shao2024privacylens developed ConfAIde and PrivacyLens, respectively, which are datasets that evaluate whether LLMs can reason about sharing data.
@@
 More recently, li2025privaci developed a benchmark targeting legal compliance.
-These datasets encompass privacy norms sourced from either legal statutes or aggregated judgments from crowdsource studies, both of which capture the general public’s privacy expectations on whether certain data sharing is appropriate in the particular context (guo2025not).
+And mireshghallah2025cimemories evaluates CI compliance of LLMs with persistent memory that was synthetically generated.
+These datasets encompass privacy norms sourced from either legal statutes or aggregated judgments from crowdsourced studies, both of which capture the general public’s privacy expectations on whether certain data sharing is appropriate in a particular context.
 Other works have operationalized the contextual awareness of LLMs to reduce privacy leakage of user tasks.
 ghalebikesabi2024operationalizing developed strategies for LLM-based assistants to be CI compliant in form filling tasks.
-bagdasarian2024airgapagent implemented a minimization module with an LLM to generate a minimized user data set based on the user task.
+bagdasarian2024airgapagent implemented a minimization module with an LLM to generate a minimized set of data based on the user’s task.
 ngong2025protecting utilized LLMs to reduce private information within a user’s prompt.
@@
 Consequently, reliance on general norms alone is ineffective for developing agents tasked with making personalized privacy decisions.
-Personalized Settings
-groschupp2025can investigated personalized access control in LLMs, where the LLM must determine whether to grant data access by evaluating an incoming request against the user’s stated access-control preferences (e.g.
-”I prefer to not share data unless necessary”).
+User Privacy Preferences
+groschupp2025can and wu2025towards investigated personalized access control in LLMs, where the LLM must determine whether to grant data access by evaluating an incoming request against a user’s access-control preferences (e.g., ”I prefer not to share data unless necessary”).
 However, it might not be practical to assume that users can explicitly specify access-control preferences.
-Further, delegating the access control decision to LLMs can struggle to align with the user’s access-control preferences, especially with limitations of LLMs such hallucinations (ji2023survey; kalai2025language; zhang2023siren) and biases (gallegos2024bias).
+Further, delegating the access control decision to LLMs may result in misalignment with the user’s access-control preferences, especially with limitations of LLMs such as hallucinations (ji2023survey; kalai2025language; zhang2023siren) and biases (gallegos2024bias).
 Another work studied how well LLMs align with human preferences in household human-robot interaction contexts (sullivan2025benchmarking).
 Our work, on the other hand, focuses on a different problem setup, where we evaluate LLMs making privacy decisions on behalf of users.
-The final work evaluates how well do judgments from LLMs align with user expectations (cheng2024ci).
-However, their evaluation is quite limited as the dataset is synthetically generated with the author’s annotating the user appropriateness labels.
+The final work evaluates how well judgments from LLMs align with user expectations (cheng2024ci).
+However, their evaluation is quite limited as the dataset is synthetically generated with the authors annotating the user-appropriateness labels.
 2.3. Neurosymbolic Approaches for Reasoning
-Prior works that are methodologically similar to ours broadly fall under neurosymbolic reasoning, specifically where the LLM acts as an interface to user inputs and offloads the reasoning through rule-based/logical approaches.
+Prior works that are methodologically similar to ours broadly fall under neurosymbolic reasoning, specifically where the LLM acts as an interface to user inputs and offloads the reasoning to rule-based/logical approaches.
 Such works include using an LLM to translate natural language into a formal intermediate language, which is then offloaded to a deterministic solver to perform the logical reasoning (olausson2023linc; pan2023logic; trinh2024solving; ye2023satlm).
-While ARIEL does not employ a theorem solver, it does use few-shot prompting with LLMs to map data-sharing requests to levels in a corresponding ontology, which can then use a rule-base component to determine entailment.
+While ARIEL does not employ a theorem solver, it does use few-shot prompting with an LLM to map data-sharing requests to levels in a corresponding ontology, which are then passed through a rule-base component to determine entailment.
 3. Personalizing Privacy Decisions
-As illustrated in Figure 1, our agent setup involves a personal agent performing tasks on a user’s behalf.
+As illustrated in Figure 1, a personal agent executes tasks on a user’s behalf.
 The agent has access to the user’s data and manages data-sharing requests from third parties required to complete these tasks.
-We describe below our problem setup of how the agent decides whether it is appropriate to transmit user data.
+Below, we formalize the problem setup describing how the agent determines whether transmitting user data is appropriate.
 3.1. Problem Statement
-We address the problem of an LLM-based agent $A_{\theta}$ tasked with making personalized privacy decisions for a user $u$.
-We assume a non-adversarial setting where the requests $R_{u}$ are not maliciously designed to manipulate the agent $A_{\theta}$.
-Upon receiving a request $R_{u}$ to share data $d_{u}$, the agent must infer the user’s judgment $l_{u}$ by leveraging a knowledge base $D_{u}$:
-$A_{\theta}(D_{u},R_{u}\in\{\text{appropriate, inappropriate}\}.$
-The knowledge base $D_{u}$ represents a history of the user’s past privacy judgments $l^{i}_{u}$ on prior requests $R^{i}_{u}$:
+We consider an LLM-based agent $A_{\theta}$ that leverages persistent memory $D_{u}$ (chhikara2025mem0; xu2025mem) to make personalized privacy decisions for a user $u$.
+We assume a non-adversarial setting where the data-sharing requests $R_{u}$ are either authenticated or otherwise not maliciously designed to manipulate the agent $A_{\theta}$.
+Upon receiving a request $R_{u}$ to share data $d_{u}$, the agent must infer the user’s judgment $l_{u}$:
+$A_{\theta}(D_{u},R_{u})\in\{\text{appropriate, inappropriate}\}.$
+The memory $D_{u}$ represents a history of the user’s past privacy judgments $l^{i}_{u}$ on data-sharing requests $R^{i}_{u}$:
 (1) $D_{u}=\{P^{i}_{u}\}_{i=1}^{n},\quad\text{where }P^{i}_{u}=(R^{i}_{u},l^{i}_{u}).$
@@
 Example 3.0.
-Let the incoming request $R_{u}$ be a bank asking for the user’s partial Social Security Number (SSN) to open a checking account.
-The knowledge base $D_{u}$ contains a prior request $R^{i}_{u}$ where the same bank asked for the user’s full SSN for the same purpose, which the user judged as appropriate ($l^{i}_{u}=\texttt{appropriate}$).
-The agent must decide if it can infer the appropriateness of $R_{u}$ based on this precedent $(R^{i}_{u},l^{i}_{u})$.
+Let the incoming request $R_{u}$ be from a bank asking for the user’s partial Social Security Number (SSN) to open a checking account.
+The memory $D_{u}$ contains a prior request $R^{i}_{u}$ in which the same bank asked for the user’s full SSN for the same purpose, which the user judged as appropriate ($l^{i}_{u}=\texttt{appropriate}$).
+The agent must decide whether it can infer the appropriateness of $R_{u}$ based on this precedent $(R^{i}_{u},l^{i}_{u})$.
 We represent each data request as a CI information flow (nissenbaum2004privacy; nissenbaum2009privacy) containing five parameters: (1) data type, the data that is being transmitted; (2) data subject, the owner of the data; (3) data sender, the individual sharing the data; (4) data recipient, the individual receiving the data; (5) transmission principle, the purpose and/or condition under which the data is shared.
@@
 Define $R_{u}=\{d_{u},u,s,r,t\}$ as a set containing five parameters where $d_{u}$ is the user’s data, $u$ is the data subject, $s$ is the data sender, $r$ is the data recipient, and $t$ is the transmission principle.
-Using Definition 3.2, we can map the prior and incoming request from Example 3.1 to the corresponding five parameters, as the following:
+Using Definition 3.2, we can map the prior and incoming requests from Example 3.1 to the corresponding five parameters, shown in Figure 2.
 Prior Request data type: full SSN
@@
 judgment:
+Figure 2. A example mapping of the prior request with user judgment and the incoming request from Example 3.1 to the five parameters.
 3.2. Baseline Approaches
 Using an LLM to address the privacy decision problem defined above requires addressing a key challenge: how to design the prompt for the LLM.
-Here, we discuss two baseline approaches to prompt the LLMs to make privacy judgments along with their limitations.
-Norm-based Baseline:
+Here, we discuss two baseline approaches to prompt the LLM to make privacy judgments, along with their limitations.
+Privacy Norms-based Baseline.
 The first choice for prompting the LLM is to utilize Privacy Norms.
 In this approach, the prompt contains three parts.
-The first part is the instruction to the LLM to make privacy decisions, the second part is the incoming request, and the third part includes the privacy norms.
-These privacy norms, akin to CI-inspired approaches, can include the general population’s overall acceptability on particular action (fan2024goldcoin).
-For instance, determining if a user should share a partial SSN with a bank would rely on broad societal acceptability.
-However, such an approach is limited because general norms lack the granularity required to capture specific, personalized user privacy decisions.
-Privacy Preferences-based Baseline:
+The first part is the instruction to the LLM to make privacy decisions, the second part includes the privacy norms, and the third part is the incoming request.
+These privacy norms, akin to CI-inspired approaches, can include the general population’s overall acceptability of a particular action (fan2024goldcoin).
+For instance, determining whether a user should share a partial SSN with a bank would rely on broad societal acceptability.
+However, such an approach is limited because general norms lack the granularity required to capture user-specific, personalized privacy decisions.
+Privacy Preferences-based Baseline.
 To address this limitation, an alternative approach is to prompt the LLM explicitly with the user’s privacy preferences.
-In this approach, the prompt also includes three parts: the general task instruction, the incoming request, and the user’s privacy preferences.
-These preferences can be based on rules or policies that the user specify, similar to groschupp2025can, or they can include a list of the user’s privacy judgments on prior requests (e.g., leveraging a past decision about sharing a full SSN to infer judgment of incoming request).
+In this approach, the prompt also includes three parts: the general task instruction, the user’s privacy preferences, and the incoming request.
+These preferences can be based on rules or policies that the user specifies, similar to groschupp2025can, or they can include a list of the user’s privacy judgments on prior requests (e.g., leveraging a past decision about sharing a full SSN to infer the judgment of incoming request) (wijesekera2017feasibility; colnago2022concern).
 Limitations
@@
 Appropriateness judgments generated directly by LLMs remain unreliable due to limited logical reasoning capabilities (lin2025zebralogic) and the potential for hallucinations that deviate from system prompts (ji2023survey; kalai2025language; zhang2023siren).
-Consequently, these judgments are not guaranteed to align with user privacy preferences, resulting in erroneous judgments that limit effective personalization.
+Consequently, these judgments are not guaranteed to align with user privacy preferences, resulting in errors that limit effective personalization.
 Furthermore, the decision-making process is opaque.
@@
 Finally, because these methods do not align with user preferences and are not inherently interpretable, they hinder user agency.
-While personalized agents should aim to mitigate privacy fatigue, this objective must not come at the cost of compromising user awareness of and control over how decisions are made on their behalf (mitchell2025fully).
+While personalized agents should aim to mitigate privacy fatigue, this objective must not come at the cost of compromising user awareness of and control over how decisions are made on their behalf (mitchell2025fully; zhang2025autonomy).
 3.3. Requirements
@@
     Alignment.
-The privacy decisions made by the agent on behalf of the users need to accurately reflect the user’s prior privacy decisions.
+The privacy decisions made by the agent on behalf of the user need to accurately reflect the user’s prior privacy decisions.
 Alignment is crucial for effective personalization of agents.
@@
     Interpretability.
-The agent needs to provide a reasoning that is transparent and traceable so that the user stays in the loop throughout the entire privacy-decision process.
+The agent needs to provide reasoning that is transparent and traceable so that the user stays in the loop throughout the privacy-decision process.
 Interpretability maintains user trust in the system.
@@
 In scenarios where the agent lacks enough information from the user’s prior privacy decisions to make a judgment, the agent must determine when it should escalate the request to the user for assistance.
-Importantly, the agent needs to accomplish the right tradeoff of automated privacy decision vs user control.
+Importantly, the agent needs to strike the right trade-off between automated privacy decisions and user control.
 4. ARIEL: Privacy Judgments with Entailment
-We propose ARIEL, a system to enable more reliable and structured reasoning for personalized privacy judgments. Below, we describe ARIEL in more detail.
+We propose ARIEL, a framework enabling aligned, structured, and interpretable reasoning for personalized privacy judgments.
+ARIEL operates as a specialized module for privacy decisions within an end-to-end agentic system (see Figure 1).
+We describe the framework in detail below.
 Figure 3.
 High-level overview of ARIEL, which can be broken down into three components.
-① ARIEL generates an ontology for each user based on their prior privacy judgments from the knowledge base using the LLM.
-② Once ARIEL receives an incoming request, it goes through each prior request in the user’s knowledge base to determine if entailment holds.
-It uses the LLM and the generated ontologies to map the parameters in both requests to correspond levels in the ontologies.
-③ A set of rules are applied to determine if entailment holds between each mapped prior requests and the mapped prior request.
+① ARIEL uses an LLM to generate an ontology for each user based on the user’s prior privacy judgments on data-sharing requests.
+② Once ARIEL receives an incoming request, it goes through each prior request from memory to determine if entailment holds.
+The LLM maps the parameters in both requests to corresponding levels in the generated ontologies.
+③ A set of rules is applied to determine whether entailment holds between each mapped prior request and the mapped incoming request.
 If the incoming request is not entailed by any prior requests, ARIEL escalates the incoming request to the user.
@@
 This approach addresses the limitations of LLM-based baselines, which do not provide a logical way to infer user judgments from prior decisions.
-In contrast, entailment treats prior judgments as logical premises that constrain future decisions, offering an interpretable and controllable method to making privacy decisions.
-This example highlights how entailment enables privacy decision based on prior requests.
+In contrast, entailment treats prior judgments as logical premises that constrain future decisions, offering an interpretable and controllable method for making privacy decisions.
+The following example highlights how entailment enables privacy decisions based on prior requests.
 Example 4.0.
@@
   1. (1)
-    Ontology Generation: use an LLM to generate relevant ontologies for each user over the parameters on incoming and prior requests.
+    Ontology Generation: use an LLM to generate relevant ontologies for each user over the parameters in each prior request and incoming request.
   2. (2)
@@
 The three components of ARIEL operate in two phases.
-In the offline phase, Ontology Generation creates ontologies for each user based on the user’s prior privacy judgments and general privacy norms encoded in the LLM.
-The online phase processes each incoming request: first, Ontology Mapping selects relevant prior requests with user judgments from the user’s knowledge base, then maps the parameters in each request to the generated ontologies.
+In the offline phase, Ontology Generation creates ontologies for each user based on the user’s prior privacy judgments and the LLM’s societal norm awareness.
+The online phase processes each incoming request: first, Ontology Mapping selects relevant prior requests with user judgments from memory, then maps the parameters in each request to the generated ontologies.
 This step provides a unified representation of the parameters in both the incoming and prior requests, enabling the entailment step.
-Next, Entailment Logic applies a set of rules to determine if entailment holds between each mapped prior requests and the mapped incoming request.
+Next, Entailment Logic applies a set of rules to determine if entailment holds between each mapped prior request and the mapped incoming request.
 If the incoming request is not entailed by any prior requests, ARIEL escalates the incoming request to the user.
@@
 ARIEL instructs the LLM to generate an ontology for each request parameter, excluding the data subject, which is always the same for each request.
-The LLM generates the ontology based on a hierarchical relationship.
+The LLM generates each ontology based on a hierarchical relationship.
 We detail each ontology with its corresponding relationship below (please refer to Appendix A.5 for the exact ontology generation prompts):
@@
   3. (3)
-    Data Recipient. Represents the user’s trust in the recipient and the recipient’s authority. Trust and authority decrease at higher levels (Level 1 has highest trust/authority).
+    Data Recipient. Represents the user’s trust in and the authority of the recipient. Trust and authority decrease at higher levels (Level 1 has highest trust/authority).
   4. (4)
-    Transmission Principle. Represents the transmission purpose and safeguards. Safeguard strength and purpose relevance decrease at higher levels (Level 1 has strongest safeguards/relevance).
+    Transmission Principle. Represents the transmission purpose and safeguards. Safeguard strength and purpose relevance decrease at higher levels (Level 1 has the strongest safeguards/relevance).
 Each hierarchical relationship is derived from two sources: (1) User’s prior privacy judgments: Relationships are inferred directly from the user’s prior privacy judgments ($D_{u}$).
@@
 This ensures the ontology is personalized.
-(2) General norms: For relationships not present in $D_{u}$, ARIEL leverages the LLM’s encoded societal norms.
+(2) General norms: For relationships not present in $D_{u}$, ARIEL leverages the LLM’s societal norm awareness.
 For example, the LLM can infer that a typical user generally trusts family more than acquaintances.
-Figure 4 illustrates a data type ontology generated by Gemini 2.5 Pro using this method.
+Figure 4 illustrates a data type ontology generated by Gemini 2.5 Pro for a particular user in the evaluation dataset using this method.
 data type Ontology L1. Non-personal, publicly available information.
@@
 For example, sharing full SSN with a bank to open a checking account is an entirely different context than sharing full SSN with an IRS agent to file taxes.
-Furthermore, even when requests are similar, the agent requires a structured process for comparing differing parameter values based on some hierarchal relationship in order to determine the entailment (e.g., comparing data recipients of parents vs friends).
+Furthermore, even when requests are similar, the agent requires a structured process for comparing differing parameter values based on some hierarchical relationship in order to determine the entailment (e.g., comparing data recipients of parents vs friends).
 To address these challenges, ARIEL employs a structured approach leveraging the generated ontologies $O$, as detailed in Algorithm 1:
-Neighboring Requests, To ensure that comparison is done on contextually similar requests, we first select prior requests that ’neighbor’ the incoming request, denoting the neighboring set as $E$.
-More precisely, $E$ contains prior requests $R^{i}_{u}$ that have a Hamming distance $d$ of at most one with the incoming request $R_{u}$.
+Neighboring Requests.
+To ensure that comparisons are made between contextually similar requests, we first select prior requests that ’neighbor’ the incoming request, denoting the neighboring set as $E$.
+More precisely, $E$ contains prior requests $R^{i}_{u}$ whose Hamming distance $d$ from the incoming request $R_{u}$ is at most one.
 This is equivalent to ensuring that both requests differ by only one parameter (line 6 in Algorithm 1).
@@
 To rigorously compare the differing parameters, we introduce an operator $\leq_{O}$ to formalize the relationship between the parameters of two distinct requests, $R$ and $R^{\prime}$, based on their mapped levels within $O$.
-Specifically, $R^{\prime}\leq_{O}R$ holds if every parameter in $R^{\prime}$ is ontologically ’less than or equal to’ the corresponding parameter in $R$.
+Specifically, $R^{\prime}\leq_{O}R$ holds if every parameter in $R^{\prime}$ is ontologically ”less than or equal to” the corresponding parameter in $R$.
 We formally define this as:
@@
 Let $R=\{d_{u},u,s,r,t\}$, $R^{\prime}=\{d^{\prime}_{u},u,s^{\prime},r^{\prime},t^{\prime}\}$ be two requests defined in Definition 3.2 that are partially ordered by the set of ontologies $O$.
-Then $R^{\prime}\leq_{O}R$ is true if $\forall p\in R,\forall p^{\prime}\in R^{\prime}$ either $p^{\prime}<_{o}p$ or $p^{\prime}=_{o}p$ for the relevant ontology $o$.
+Then $R^{\prime}\leq_{O}R$ holds if, for each corresponding parameter $p\in R$ and $p^{\prime}\in R^{\prime}$, either $p^{\prime}<_{o}p$ or $p^{\prime}=_{o}p$ under the relevant ontology $o\in O$.
 Then ARIEL iterates through each neighboring prior request $R^{\prime}_{u}$.
-For each pair, the agent retrieves the differing parameter values, $p$ (from $R$) and $p^{\prime}$ (from $R^{\prime}_{u}$).
+For each pair, the agent retrieves the differing parameter values, $p$ (from $R_{u}$) and $p^{\prime}$ (from $R^{\prime}_{u}$).
 These values are then mapped to the relevant ontology $o\in O$ (line 12 in Algorithm 1).
-For this ontological mapping, we use the LLM $\theta$ to determine the relevant ontology $o$ and map $p$ and $p^{\prime}$ to their corresponding levels in $o$
+For this ontological mapping, we use the LLM $\theta$ to determine the relevant ontology $o$ and map $p$ and $p^{\prime}$ to their corresponding levels in $o$.
 4.4. Entailment Logic
@@
 Let $R,R^{\prime}$ be two requests with $l^{\prime}$ being an appropriateness label for $R^{\prime}$.
-And let $O$ be a set of ontologies.
-We say that the entailment $(R^{\prime},l^{\prime})\implies(R,l^{\prime})$ holds if one of the following two cases hold:
+Let $O$ be a set of ontologies.
+We say that the entailment $(R^{\prime},l^{\prime})\implies(R,l^{\prime})$ holds if one of the following two cases holds:
   1. (1)
@@
 To address this, the agent goes through each neighboring request and keeps a running vote of the appropriate and inappropriate judgments that were entailed.
-Whichever appropriateness receives the majority vote is the final judgment.
-If there is no majority, then the agent outputs ’undetermined’ (line 19 in Algorithm 1), signifying that the incoming request should be escalated to the user for further assistance.
-Lastly, we note that in scenarios where multiple appropriate and inappropriate prior requests are entailed, ARIEL can be modified for the agent to escalate the incoming request to the user instead of employing majority vote.
+Whichever appropriateness label receives the majority vote is the final judgment.
+If there is no majority, the agent outputs ’undetermined’ (line 19 in Algorithm 1), signifying that the incoming request should be escalated to the user for further assistance.
+Lastly, we note that in scenarios where multiple appropriate and inappropriate prior requests are entailed, ARIEL can be modified so that the agent to escalate the incoming request to the user instead of employing a majority vote.
 5. Evaluation Setup
@@
 Participants are presented with data-sharing scenarios and asked to indicate their acceptance of each one (abdi2021privacy; apthorpe2018discovering; apthorpe2019evaluating; bhatia2018empirical; hoyle2020privacy; martin2016measuring; shvartzshnaider2016learning).
-To simulate a user with prior privacy judgments, each of the data-sharing scenarios are converted into requests with appropriateness judgments from the participants.
+To simulate a user with prior privacy judgments, each data-sharing scenario is converted into a request (Definition 3.2) with appropriateness judgments directly received from the participants.
 Then, we divided a participant’s set of answered requests into two pairwise disjoint subsets.
@@
 The second subset represents incoming requests where the participant’s answers were held out as the ground truth for evaluation.
-Our evaluation was able to employ two such datasets, which we repurposed for our setup as detailed below.
+Our evaluation employs two such datasets, which we repurposed for our setup as detailed below.
 See Figure 5 for example requests from both datasets.
@@
 The scenarios typically involve a voice assistant provider sending user data to individuals/entities who have access to the user’s SPA ecosystem, under a certain purpose and condition.
-The data subject of a request is always ’user’ and the data sender is always ’assistant provider’.
+The data subject of a request is always ”user” and the data sender is always ”assistant provider”.
 The total number of participants is 1,730, with each participant randomly assigned to answer a subset of the questions.
@@
 For each of the 500 users, we partitioned the responses into 60 prior requests and 10 incoming requests.
-This yields a total of 5,000 incoming requests, consisting of 1,765 appropriate and 3,235 inappropriate labels.
+This yields a total of 50010=5000500*10=5000 incoming requests, consisting of 1,765 appropriate and 3,235 inappropriate labels.
 We also perform an ablation study on varying the number of prior requests in Section 6.2.
-Lastly, in the dataset, the participant responses are in 1-5 Likert scale where 1-2 correspond to strongly and somewhat unacceptable, 3 corresponds to neutral, and 4-5 correspond to somewhat and strongly acceptable.
+Lastly, in the dataset, the participant responses are on a 1-5 Likert scale where 1 and 2 correspond to strongly and somewhat unacceptable responses, respectively, 3 corresponds to neutral, and 4 and 5 correspond to somewhat and strongly acceptable responses, respectively.
 We convert the responses with 1-2 scores to ”Inappropriate” and the responses with 4-5 to ”Appropriate”.
@@
 The participant responses in the dataset contain numerical values, where 1 corresponds to Yes, 2 corresponds to No, and 3-5 correspond to Does not Make Sense options.
-Specifically, 3 is ’The sender is unlikely to have the information, 2 is ’The receiver would already have the information, and 3 is ’the question is ambiguous.’ We convert responses with 1 to ”Appropriate” and with 2 to ”Inappropriate”.
+Specifically, 3 is ”The sender is unlikely to have the information”, 4 is ”The receiver would already have the information”, and 5 is ”the question is ambiguous.” We convert responses with 1 to ”Appropriate” and with 2 to ”Inappropriate”.
 We filter out the remaining responses with 3-5 scores.
-Lastly, we filter out users that have less than 70 appropriateness judgments, so each user has 60 prior requests and 10 incoming requests.
+Lastly, we filter out users who have fewer than 70 appropriateness judgments, so each user has 60 prior requests and 10 incoming requests.
 In total, there are 302 users and 30210=3020302*10=3020 total incoming requests, consisting of 1378 appropriate labels and 1642 inappropriate labels.
@@
 Following the description in Section 3.2, we employ two natural baseline prompt strategies: (1) privacy norms representative of the general population and (2) prior privacy judgments from users.
-5.2.1. Norms-based Baseline
-Privacy norms capture the general population’s accepted appropriateness of information flows. There are two prompts that we consider.
+5.2.1. Privacy Norms-based Baseline
+Privacy norms capture the general population’s accepted appropriateness of information flows. We consider two prompts.
   1. (1)
@@
 The prompt directs the LLM to judge the appropriateness of each information flow based on its knowledge of general privacy norms.
-Prompts can be found in Appendix A.1.
+The prompts can be found in Appendix A.1.
   2. (2)
@@
 We construct the prompt such that the LLM’s privacy judgment is now based on a provided subset of representative privacy norms.
-These extracted privacy norms are taken directly from the analysis provided in the papers of each dataset.
-Table 2 contains a subset of the general privacy from both evaluation datasets that were used by the LLMs.
+These extracted privacy norms are taken directly from the analysis provided in the original papers for each dataset.
+Table 2 contains a subset of the general privacy norms from both evaluation datasets that were used by the LLMs.
 The prompt directs the LLM to judge the appropriateness of each information flow based on the provided general privacy norms.
-Prompts can be found in A.2.
+The prompts containing the entire set of privacy norms can be found in Appendix A.2.
 Rule Judgment
@@
 5.2.2. Privacy Preferences-based Baseline
-We provide privacy judgments on prior requests made by a user. There are two prompts that we then consider.
+We provide privacy judgments on prior requests made by a user. We consider two prompts.
   1. (1)
@@
 We modify the instruction in ICL to output ’undetermined’ if the model is unable to confidently judge whether the incoming request is appropriate or inappropriate based on the user’s privacy preferences.
-These account for scenarios where the user’s privacy preferences may not always provide sufficient information to accurately judge the incoming request.
+This modification accounts for scenarios where the user’s privacy preferences may not always provide sufficient information to accurately judge the incoming request.
 The prompts are in A.4.
+5.3. Models and Evaluation metrics
+Models
+We evaluate ARIEL and the baselines on three advanced models: Gemini 2.5-Pro (gemini-2.5-pro), GPT-5 (gpt-5-2025-08-07), and Claude Sonnet 4.5 (claude-sonnet-4-5-20250929).
+Additionally, we include a smaller, open-source model, Gemma 3 4B IT (team2025gemma).
+Metrics
+We compare the agent’s predicted judgments against each sampled user’s judgments on incoming requests.
+We report separate F1 scores for both the appropriate and inappropriate classes.
+Given an evaluation dataset $D$ containing incoming requests $R_{u}$ with judgments $l_{u}$ for each user $u$, we define the precision $P_{A}$ for the appropriate class ($c^{+}=\texttt{appropriate}$) as:
+$P_{A}:=\frac{|(R_{u},l_{u})\in D:A_{\theta}(D_{u},R_{u})=c^{+}\wedge l_{u}=c^{+}|}{|(R_{u},l_{u})\in D:A_{\theta}(D_{u},R_{u})=c^{+}|}.$
+We also define recall $R_{A}$ for the appropriate class as:
+$R_{A}:=\frac{|(R_{u},l_{u})\in D:A_{\theta}(D_{u},R_{u})=c^{+}\wedge l_{u}=c^{+}|}{|(R_{u},l_{u})\in D:l_{u}=c^{+}|}.$
+Then we can define the F1 score $F1_{A}$ for the appropriate class as:
+$F1_{A}=2\cdot\frac{P_{A}\cdot R_{A}}{P_{A}+R_{A}}.$
+A similar process can be used to obtain the F1 score for the inappropriate class $F1_{I}$.
+High $F1_{A}$ indicates the agent is correctly sharing user data (utility), whereas high $F1_{I}$ indicates the agent is correctly withholding data (privacy).
+Evaluation Procedure for general norms.
+For each unique request in the datasets, we generate a prompt as described in Section 5.2.1.
+This results in 825 prompts for the SPA dataset and 1411 for the Education dataset, covering both zero-shot and privacy norms.
+Once the prompts are generated, we provide all of them to the models to obtain their responses.
+Then, for each user, we compare their judgment on an incoming request with the LLM’s judgment on that same request across both general norms prompt types.
+Evaluation Procedure for user privacy preferences.
+We generate a prompt for each incoming request with the user’s privacy judgments on prior requests.
+This results in 5000 prompts for the SPA dataset and 3020 prompts for the Education dataset, covering ICL, ICL (w Undet), and ARIEL.
+Incoming requests resulting in an ’undetermined’ judgment are excluded from the evaluation for the respective baselines (i.e., ICL (w Undet) or ARIEL).
+To account for this partial coverage, we report the Support (Sup) metric, defined as the total number of appropriateness judgments (i.e., ”appropriate” or ”inappropriate”) made by the model.
+6. Results
 Dataset Model zero-shot privacy norms ICL ICL (w Undet) ARIEL
@@
 Claude Sonnet 4.5 40.0 75.0 57.7 73.6 81.4 91.0 5000 83.3 91.7 4900 86.6 93.8 3824
-Gemma 3 4B 45.1 72.8 46.4 73.1 64.1 84.4 5000 62.5 84.2 4999 75.5 88.4 3680
+Gemma 3 4B IT 45.1 72.8 46.4 73.1 64.1 84.4 5000 62.5 84.2 4999 75.5 88.4 3680
 Education Gemini 2.5 Pro 60.1 61.7 38.5 71.6 69.9 80.8 3020 75.5 87.2 2325 84.1 88.2 1681
@@
 Claude Sonnet 4.5 61.4 63.0 62.3 62.5 75.4 82.8 3020 75.0 82.8 2950 84.0 88.4 1608
-Gemma 3 4B 64.1 18.0 46.7 66.8 61.7 78.7 3020 63.4 78.8 3020 80.4 86.1 1644
+Gemma 3 4B IT 64.1 18.0 46.7 66.8 61.7 78.7 3020 63.4 78.8 3020 80.4 86.1 1644
 Table 3.
@@
 These results suggest that ARIEL accurately identifies when the appropriateness of incoming requests are difficult to judge based on the user’s prior privacy judgments.
-5.3. Models and Evaluation metrics
-Models
-We evaluate ARIEL and the baselines on three advanced models: Gemini 2.5-Pro (gemini-2.5-pro), GPT-5 (gpt-5-2025-08-07), and Claude Sonnet 4.5 (claude-sonnet-4-5-20250929).
-Additionally, we include a smaller, open source model, Gemma 3 4B (team2025gemma).
-Metrics
-We compare the agent’s predicted judgments against the sampled users’ judgments on incoming requests.
-We report separate F1 scores for both the appropriate and inappropriate classes.
-Given an evaluation dataset $D$ containing incoming requests $R_{u}$ with judgments $l_{u}$ for each user $u$, we define the precision ($P_{A}$) for the appropriate class ($c^{+}=\texttt{appropriate}$) as:
-$P_{A}:=\frac{|(R_{u},l_{u})\in D:A_{\theta}(D_{u},R_{u})=c^{+}\wedge l_{u}=c^{+}|}{|(R_{u},l_{u})\in D:A_{\theta}(D_{u},R_{u})=c^{+}|}.$
-We also define recall ($R_{A})$ for the appropriate class as:
-$R_{A}:=\frac{|(R_{u},l_{u})\in D:A_{\theta}(D_{u},R_{u})=c^{+}\wedge l_{u}=c^{+}|}{|(R_{u},l_{u})\in D:l_{u}=c^{+}|}.$
-Then we can define the F1 score ($F1_{A}$) for the appropriate class as:
-$F1_{A}=2\cdot\frac{P_{A}\cdot R_{A}}{P_{A}+R_{A}}.$
-A similar process can be done to obtain the F1 Score for the inappropriate class $F1_{I}$.
-Evaluation Procedure for general norms.
-For each unique request in the datasets, we generate a prompt as described in Section 5.2.1.
-This results in 825 prompts for the SPA dataset and 1,411 for the Education dataset, covering both zero-shot and privacy norms.
-Once the prompts are generated, we provide all of them to the models to obtain their responses.
-Then, for each user, we compare their judgment on an incoming request with the LLM’s judgment on that same request, evaluating this on both general norms prompt types.
-Evaluation Procedure for user privacy preferences.
-We generate a prompt for each incoming request with the user’s privacy judgments on prior requests.
-This results in 5,000 prompts for the SPA dataset and 3020 prompts for the Education dataset, covering ICL, ICL (w Undet), and ARIEL.
-Incoming requests resulting in an ’undetermined’ judgment are excluded from the evaluation for the respective baselines (i.e., ICL (w Undet) or ARIEL).
-To account for this partial coverage, we report the Support (Sup) metric, defined as the total number of appropriateness judgments (i.e., ’appropriate’ or ’inappropriate’) made by the model.
-6. Results
 We present experiments, ablation studies, and error analysis to study the efficacy and robustness of ARIEL.
-First we evaluate the end-to-end performance improvement of ARIEL over existing baselines in Section 6.1.
-Then we evaluate how varying the number of prior user judgments affect ARIEL in Section 6.2.
+First, we evaluate the end-to-end performance improvement of ARIEL over existing baselines in Section 6.1.
+Then, we evaluate how varying the number of prior user judgments affects ARIEL in Section 6.2.
 Finally, we qualitatively analyze the failure modes of ARIEL in Section 6.3.
@@
 For example, ARIEL improves upon ICL (w Undet) by 10.4% $F1_{\text{A}}$ with GPT-5 on the Education dataset.
-In other words, the $F1_{\text{A}}$ error reduces by 39.1%, demonstrating that logical entailment effectively grounds judgments in prior user privacy judgments.
+In other words, the $F1_{\text{A}}$ error reduces by 40.6%, demonstrating that logical entailment effectively grounds judgments in prior user privacy judgments.
 Notably, these improvements are magnified in smaller models.
-With Gemma 3 4B, ARIEL yields $F1_{\text{A}}$ gains of 13% and 17% over ICL (w Undet) on the SPA and Education datasets, respectively.
+With Gemma 3 4B IT, ARIEL yields $F1_{\text{A}}$ gains of 13% and 17% over ICL (w Undet) on the SPA and Education datasets, respectively.
 This demonstrates ARIEL is suitable for on-device deployment of personalized agents.
-Figure 6.
-Ablation study on the number of prior requests with user judgments contained in the user’s knowledge base $D_{u}$.
-We evaluate ICL w (Undet) and ARIEL on the SPA dataset with Gemini 2.5 Pro.
-We report the F1-score for appropriate $F1_{\text{A}}$ and inappropriate $F1_{\text{I}}$ class, and the total number of appropriateness judgments (Sup).
-We find that ARIEL is more robust to varying the number of prior requests compared to ICL w (Undet)
-Figure 7.
-Ablation study on CoT.
-We evaluate zero-shot, privacy norms, and ICL w (Undet) on the SPA dataset with Gemini 2.5 Pro.
-We report the F1-score for appropriate $F1_{\text{A}}$ and inappropriate $F1_{\text{I}}$ class.
-The results suggest that generally CoT has negligible impact on privacy judgments across both models.
 (2) ARIEL accurately identifies when the appropriateness of an incoming request is challenging to judge.
-We observe that the number judgments that were undetermined by ARIEL is higher than ICL (w Undet).
-For Gemini 2.5 Pro, we see that ARIEL judged 1203 and 1339 incoming requests as undetermined for SPA and Education datasets, respectively, compared to just 387 and 695 for ICL (w Undet).
+We observe that the number of judgments that were undetermined by ARIEL is higher than ICL (w Undet).
+For Gemini 2.5 Pro, we observe that ARIEL judged 1203 and 1339 incoming requests as undetermined for SPA and Education datasets, respectively, compared to just 387 and 695 for ICL (w Undet).
 This observation raises the question of whether the appropriateness label of the incoming requests left undetermined by ARIEL can actually be inferred by an LLM based on the user’s prior privacy judgments.
@@
 We observe that performance on appropriate judgments from zero-shot is low for both models and datasets.
-For inappropriate judgments, zero-shot prompting does performs much better than on appropriate judgments, but the results remain suboptimal.
+For inappropriate judgments, zero-shot prompting can perform much better than on appropriate judgments, but the results remain suboptimal.
 Hence, the LLM’s awareness of privacy norms cannot sufficiently capture the nuanced behavior of user privacy preferences within the SPA and Education contexts.
-(4) General privacy norms extracted from overall user responses for each dataset can improve performance over zero-shot.
-We observe that privacy norms improve appropriate judgments for the SPA dataset across all models, and inappropriate judgments for the Education dataset on most.
+(4) General privacy norms extracted from overall user responses for each dataset can improve personalized privacy judgments.
+When comparing privacy norms and zero-shot, we observe that providing privacy norms to the LLMs can improve appropriate judgments for the SPA dataset across all models, and improve inappropriate judgments for the Education dataset on most models.
 However, including general norms can sometimes degrade performance.
-For instance, with Gemini 2.5 Pro on the Education dataset, $F1_{\text{A}}$ decreases by $10\%$ while $F1_{\text{I}}$ improves by only $8\%$, suggesting that adding general privacy norms is not always beneficial.
+For instance, with Gemini 2.5 Pro on the Education dataset, $F1_{\text{A}}$ decreases by $22\%$ while $F1_{\text{I}}$ improves by only $10\%$, suggesting that adding general privacy norms is not always beneficial.
+(5) Prior user privacy judgments improve personalized privacy judgments over general privacy norms.
+By comparing the $F1_{\text{A}}$ and $F1_{\text{I}}$ between ICL and both zero-shot and privacy norms for all datasets and models, it is clear that personalizing LLMs with prior user privacy judgments improves appropriate and inappropriate judgments over general privacy norms.
+In some cases, the improvement over general privacy norms can be as large as 50% for $F1_{\text{A}}$.
+(6) Personalized privacy judgments improve when including undetermined as an option.
+When comparing ICL and ICL (w Undet), we observe an improvement in $F1_{\text{A}}$ and $F1_{\text{I}}$ for all LLMs.
+These results suggest that giving an LLM the option to abstain reduces the likelihood that the model forces an erroneous appropriateness judgment.
+6.2. Ablation Evaluation
+We evaluate how the number of prior requests affects the performance of ARIEL and baseline approaches. Moreover, we investigate how the choice of reasoning affects LLM privacy judgments.
+Figure 6.
+Ablation study on the number of prior requests with user judgments contained in user memory $D_{u}$.
+We evaluate ICL (w Undet) and ARIEL on the SPA dataset with Gemini 2.5 Pro.
+We report the F1-score for appropriate $F1_{\text{A}}$ and inappropriate $F1_{\text{I}}$ class, and the total number of appropriateness judgments (Sup).
+We find that ARIEL is more robust to varying the number of prior requests compared to ICL w (Undet).
+Figure 7.
+Ablation study on CoT.
+We evaluate zero-shot, privacy norms, and ICL w (Undet) on the SPA dataset with Gemini 2.5 Pro.
+We report the F1-score for the appropriate $F1_{\text{A}}$ and inappropriate $F1_{\text{I}}$ class.
+The results suggest that generally CoT has negligible impact on privacy judgments across both models.
+6.2.1. Number of prior requests
+We explore how varying the number of prior requests with user judgments affects the agent’s judgments.
+Specifically, we compare the $F1_{A}$, $F1_{I}$, and Support (Sup) of ARIEL against ICL (w Undet) on the SPA dataset using Gemini 2.5 Pro.
+As evident in Figure 6, while the $F1_{I}$ gap remains relatively constant, the disparity in $F1_{A}$ widens significantly as the number of prior requests decreases.
+Notably, when reducing the prior requests from 60 to 20, $F1_{A}$ drops by only 3.4% for ARIEL, compared to an 8.8% decline for ICL (w Undet).
+This demonstrates that ARIEL is more robust to limited prior privacy judgments.
+Nonetheless, we observe that the number of entailments generated by ARIEL decreases substantially as the number of prior requests decreases.
+This result is expected, as there are fewer candidate requests available to satisfy the entailment criteria (i.e., fewer prior requests differ by exactly one parameter from the incoming request).
+6.2.2. Assessing the impact of reasoning with Chain of Thought (CoT)
+To assess the impact of reasoning strategy choice on appropriateness judgments, we implement CoT by prompting the LLM to provide a reasoning before making a judgment (wei2022chain; yang2023large).
+Inspired by findings that CoT affects privacy leakage (mireshghallah2024can), we evaluate its effect on judgment accuracy across zero-shot, privacy norms, and ICL (w Undet).
+Figure 7 indicates that CoT typically provides negligible changes in performance.
+However, we observe for zero-shot with Gemini 2.5 Pro that the $F1_{A}(\%)$ and $F1_{I}(\%)$ are significantly impacted by CoT, both positively and negatively.
+While CoT provides mixed results for general privacy norms, the results clearly demonstrate that adding CoT to both ICL and ICL (w Undet) consistently hurts their performance.
+When manually inspecting the generated reasoning traces, we notice that the LLM makes misaligned assertions about the user’s privacy judgments, which propagate all the way to the model’s final judgment (see Section 6.3 for more information).
+This result also demonstrates that eliciting explanations from LLM-based reasoning can degrade appropriateness judgments.
+In contrast, ARIEL provides a clear, traceable reasoning while maintaining well-grounded personalized privacy judgments.
+6.3. Qualitative Analysis
+Finally, we qualitatively analyze incorrect judgments made by both ICL (w Undet) + CoT and ARIEL.
+We sample 20 examples from each method on the SPA dataset with Gemini 2.5 Pro.
+The results are detailed in Table 5.
+6.3.1. Error Analysis on Appropriateness Judgments for ICL (w Undet) + CoT
+Analysis of the reasoning traces reveals that Gemini 2.5 Pro occasionally performs implicit entailment when identifying prior requests that share identical parameters except for one.
+This behavior is noteworthy, as we did not explicitly instruct the model to use this logic, but only to extract patterns from user privacy judgments.
+We find that the majority of errors stem from preference misalignment (14 examples), where the model draws incorrect equivalencies between distinct parameter values to perform this entailment.
+For instance, given prior requests involving ”close friends” or ”parents” as recipients and an incoming request involving ”housemates,” the model infers that the privacy judgment should remain consistent due to perceived similarity in recipient type.
+However, the user may not view these three recipients as similar.
+We also attribute a subset of errors to the model’s inherent conservatism regarding information sharing, which we describe as default to inappropriate (5 examples).
+This occurs when the model concludes it cannot use the most similar prior requests that are judged as ”appropriate” by the user.
+However, rather than outputting ”undetermined”, the model reasons that the judgment is inappropriate.
+Finally, we identified one instance of contradictory prior requests, where the provided user history contained judgments that logically conflicted with the ground truth for the incoming request.
 Type Example/Description Total
@@
 Table 5. The different error types of 20 judgments from ICL (w Undet) + CoT and ARIEL on SPA dataset with Gemini 2.5 Pro.
-(5) Prior user privacy judgments improve personalized privacy judgments over general privacy norms.
-By comparing the $F1_{\text{A}}$ and $F1_{\text{I}}$ between ICL and both zero-shot and privacy norms for all datasets and models, it is clear that personalizing LLMs with prior user privacy judgments improves appropriate and inappropriate judgments over general privacy norms.
-The improvement over general privacy norms can in some cases be as large as 20% for $F1_{\text{A}}$.
-(6) Personalized privacy judgments improve when including undetermined as an option.
-When comparing ICL and ICL (w Undet), we observe an improvement in $F1_{\text{A}}$ and $F1_{\text{I}}$ for all state-of-art reasoning models.
-These results suggest that a hybrid approach of combining automated agent judgments with occasional escalation to the user can be highly effective for personalization.
-6.2. Ablation Evaluation
-We evaluate how the number of prior request affect the performance of ARIEL and baseline approaches. Moreover, we investigate how the choice of reasoning affects LLM privacy judgments.
-6.2.1. Number of prior requests
-We explore how varying the number of prior requests with user judgments affect the agent’s judgments.
-Specifically, we compare the $F1_{A}$, $F1_{I}$, and Support (Sup) of ARIEL against ICL (w Undet) on the SPA dataset using Gemini 2.5 Pro.
-As evident in Figure 6, while the $F1_{I}$ gap remains relatively constant, the disparity in $F1_{A}$ widens significantly as the number of prior requests decreases.
-Notably, when reducing the prior requests from 60 to 20, $F1_{A}$ drops by only 3.4% for ARIEL, compared to an 8.8% decline for ICL (w Undet).
-This demonstrates that ARIEL is more robust to limited prior privacy judgments.
-Nonetheless, we observe that the number of entailments generated by ARIEL decreases substantially as the number of prior requests decreases.
-This result is expected, as there are fewer candidate requests available to satisfy the entailment criteria (i.e., fewer prior requests differ by exactly one parameter from the incoming request).
-6.2.2. Assessing impact of reasoning with Chain of Thought (CoT)
-To assess the impact of reasoning strategy choice on appropriateness judgments, we implement CoT by prompting the LLM to provide a reasoning before making a judgment (wei2022chain).
-Inspired by findings that CoT affects privacy leakage (mireshghallah2024can), we evaluate its effect on judgment accuracy across zero-shot, privacy norms, and ICL (w Undet).
-Figure 7 indicates that CoT typically provides negligible changes in performance.
-However, we observe for zero-shot with Gemini 2.5 Pro that the $F1_{A}(\%)$ and $F1_{I}(\%)$ is significantly impacted by CoT, both positively and negatively.
-While CoT provides mixed results for general privacy norms, the results clearly demonstrate that adding CoT to both ICL and ICL (w Undet) consistently hurts their performance.
-When manually inspecting the generated reasoning traces, we notice that the LLM makes misaligned assertions about the user’s privacy judgments, which propagate all the way to the model’s final judgment (see Section 6.3 for more information).
-This result also demonstrates that eliciting explainability from LLM-based reasoning can degrade the appropriateness judgment.
-In contrast, ARIEL provides a clear, traceable reasoning while maintaining well-grounded personalized privacy judgments.
-6.3. Qualitative Analysis
-Finally, we qualitatively analyze incorrect judgments made by both ICL (w Undet) + CoT and ARIEL.
-We sample 20 examples from each method on the SPA dataset with Gemini 2.5 Pro.
-The results are detailed in Table 5.
-6.3.1. Error Analysis on incorrect appropriateness judgments for ICL (w Undet) + CoT
-Analysis of the reasoning traces reveals that Gemini 2.5 Pro occasionally performs implicit entailment when identifying prior requests that share identical parameters except for one.
-This behavior is noteworthy, as we did not explicitly instruct the model to use this logic, but only to extract patterns from user privacy judgments.
-We find that the majority of errors stem from preference misalignment (14 examples), where the model draws incorrect equivalencies between distinct parameter values to do this entailment.
-For instance, given prior requests involving ’close friends’ or ’parents’ as recipients and an incoming request involving ’housemates,’ the model infers that the privacy judgment should remain consistent due to perceived similarity in type of recipient.
-However, the user may not view these three recipients as similar.
-We also attribute a subset of errors to the model’s inherent conservatism regarding information sharing, which we describe as default to inappropriate (5 examples).
-This occurs when the model concludes it cannot use the most similar prior requests that are judged as ’appropriate’ by the user.
-But rather than outputting ’undetermined’, the model reasons that the judgment is inappropriate.
-Finally, we identified one instance of contradictory prior requests, where the provided user history contained judgments that logically conflicted with the ground truth for the incoming request.
-6.3.2. Error analysis on incorrect appropriateness judgments for ARIEL
+6.3.2. Error Analysis on Appropriateness Judgments for ARIEL
 We classify errors in ARIEL into three categories: (1) ontology generation, (2) ontology mapping, and (3) contradictory prior requests.
@@
 This is often caused by domain-specific idiosyncrasies in the evaluation dataset.
-For instance, the transmission principle has specific format combining both purpose and condition, which can be challenging for the model to map.
+For instance, the transmission principle has a specific format combining both purpose and condition, which can be challenging for the model to map.
 Finally, the remaining errors (5 examples) result from contradictory prior requests.
-For instance, there was a prior request where user had judged sharing bank account details with business and finance skills for the purpose of making transactions as inappropriate.
-The incoming request is identical as the prior, except there is no purpose.
+For instance, there was a prior request where the user had judged sharing bank account details with business and finance skills for the purpose of making transactions as inappropriate.
+The incoming request is identical to the prior, except there is no purpose.
 Thus, the incoming request should also be inappropriate, since removing the purpose likely makes the request more inappropriate.
-Both ICL (w Undet) +CoT and ARIEL judged this request inappropriate, however, the user judged the incoming request to be appropriate.
+Both ICL (w Undet) + CoT and ARIEL judged this request inappropriate; however, the user judged the incoming request to be appropriate.
+Potential solutions for (1) involve providing more user judgments on prior requests during ontology generation, while solutions for (2) involve providing more information on the domain-specific format of each parameter in the ontology mapping prompt.
+The errors in (3) represent an open challenge that is discussed in more detail in Section 7.1.
 7. Discussion
-Based on the results and takeaways, we discuss in more details the broad implications of this work as well as its limitations.
-7.1. Personalization and Privacy Norms
-Our results in Section 6 suggest that effective personalization of LLM agents requires grounding judgments in prior user privacy judgments, as general norms are insufficient.
-In this paper, we consider user judgments to prior requests as proxies for the user’s privacy preferences.
-However, this finding does not invalidate use general norms in LLMs.
-Aligning models with relevant contextual norms remains crucial for many data-sharing scenarios (mireshghallah2024can; shao2024privacylens).
-Our work highlights that when personalization is a crucial component for making privacy decisions, then the LLM’s privacy judgments should be grounded in user-specific preferences rather than general privacy norms.
+Based on our results and takeaways, we discuss in more detail the broad implications of this work as well as its limitations.
 7.1. Personalization Challenges
-Personalizing LLM-based agents introduces several challenges that must be considered.
-Contradictory privacy preferences.
-The first challenge, highlighted by the error analysis in Section 6.3, arises from contradictions within the user’s own history of prior judgments.
+In this paper, we consider the user’s judgments on prior data-sharing requests as proxies for the user’s privacy preferences (wijesekera2017feasibility; colnago2022concern).
+Personalizing privacy decisions based on these preferences introduces several challenges that must be considered.
+Contradictory Privacy Preferences
+Highlighted by the error analysis in Section 6.3, contradictions can arise within the user’s own history of prior judgments.
 For example, the user says it is inappropriate to share their banking information with a third party if their data was confidential, but deems it appropriate if there is no safeguard on the data.
@@
 Consequently, the agent must (1) detect these user privacy judgment contradictions and (2) resolve them, typically by escalating the request to the user.
-The fundamental difficulty lies in distinguishing whether an incorrect privacy judgment stems from an error in the entailment process or an inconsistency in the user’s prior privacy judgments.
-Evolving privacy preferences.
+The fundamental difficulty lies in distinguishing whether an incorrect privacy judgment stems from an error in the entailment process of ARIEL or an inconsistency in the user’s prior privacy judgments.
+Evolving Privacy Preferences
 The second challenge involves the dynamic nature of privacy preferences as they can change over time.
 For example, a user can become more conservative about sharing their data as they get older.
-Hence, a system similar to ARIEL requires a mechanism to detect when the user’s privacy preferences are shifting and determine which preferences must be updated.
-One potential mechanism could trade-off exploration and exploitation by escalating incoming requests to the user every once in a while to confirm the agent’s decision.
+Currently, ARIEL does not support automatic detection of privacy preference changes.
+Hence, future adaptations of ARIEL require a mechanism to detect when the user’s privacy preferences are shifting and determine which preferences must be updated.
+One potential mechanism is to periodically escalate incoming requests to the user to confirm the agent’s privacy decision.
 7.2. User Agency in Personal AI Agents
-While personalized agents reduce the burden of privacy management, they must not remove user agency (mitchell2025fully).
-This observation presents a tradeoff where the agent automates clear decisions but escalates cases where the judgment cannot be easily inferred by the user’s prior judgments.
-ARIEL implements this via an ’undetermined’ option, and our results confirm that identifying these non-inferable cases drives ARIEL’s judgment accuracy gains.
+While personalized agents reduce the burden of privacy management, they must not remove user agency (mitchell2025fully; zhang2025autonomy).
+This observation motivated the entailment logic of ARIEL, where the agent automates clear decisions that can be entailed but escalates requests where the judgment cannot be easily inferred by the user’s prior judgments.
+Our results confirm that identifying these non-entailed incoming requests drives ARIEL’s judgment accuracy gains.
 Furthermore, by providing transparent reasoning traces for user auditing, ARIEL successfully balances automation with necessary human oversight.
-Future work could operationalize this transparency through a dedicated user interface, enabling users to visualize the agent’s logic and intervene in decision-making when necessary.
+Although ARIEL was designed with user agency as a core requirement, further work is needed to strengthen the interaction between the user and the agent.
+One promising direction is the development of a dedicated user interface (UI) that enables users to: (1) visualize the reasoning steps behind ARIEL’s decisions (e.g., which requests were entailed and how they were mapped in the ontology), and (2) intervene in the decision-making process when necessary.
+Subsequently, this UI can facilitate user studies to evaluate how well ARIEL’s privacy reasoning and judgments align with the user’s own privacy reasoning (nezhad2026understanding).
+Furthermore, this UI could be used to evaluate potential solutions for the challenges outlined in Section 7.1, such as user verification of ARIEL’s reasoning and decisions to identify contradictions or changes in the user’s privacy preferences.
+7.3. Personalization and Privacy Norms
+Our results in Section 6 suggest that effective personalization of LLM agents requires grounding judgments in prior user privacy judgments, as general norms are insufficient.
+However, this finding does not invalidate alignment of LLMs with general contextual norms, as this alignment remains crucial for many data-sharing scenarios (mireshghallah2024can; shao2024privacylens).
+Our work highlights that when personalization is a crucial component for making privacy decisions, then the LLM’s privacy judgments should be grounded in user-specific preferences rather than general privacy norms.
 7.4. Limitations
-There are three main limitations involving our evaluation in this work.
-The first is that our evaluation datasets simulate prior requests by partitioning the users requests into two pairwise disjoint sets, where one set stores already observed requests while the other contains new requests.
+We identify four main limitations involving our work.
+Ontology Mapping and Generation
+ARIEL generates generalizable ontologies for individual fields (e.g., data type) based on the user’s prior privacy judgments.
+However, in practice, the ontology of one field may depend on the values of others.
+For instance, data sensitivity can vary based on the recipient: a user may consider their income more sensitive than SSN when sharing with a parent, but the sensitivity is reversed when sharing with a friend.
+Our results in Section 6.1 demonstrate that even without explicitly modeling these conditional dependencies, ARIEL accurately infers privacy judgments.
+We leave the incorporation of context-conditional hierarchies for future work.
+Evaluation Datasets
+Our evaluation datasets simulate prior requests by partitioning the user’s judgments on requests into two pairwise disjoint sets, where one set stores already observed requests while the other contains new requests.
 While we believe this reasonably approximates our problem setup, it does not incorporate user decision-making over time.
-Unfortunately, to our knowledge, there are no existing publicly-available datasets that incorporates user privacy preferences over time.
-An interesting future work could elicit user privacy preferences over a period of time, rather than giving a one-time questionnaire to participants.
-The second is our evaluation does not consider cold-start cases where the user has not made prior judgments on requests– essentially, no privacy preferences.
-While falling back to general privacy norms is a potential solution, our results indicate they are often insufficient for personalized privacy decisions.
-An alternative solution could be to develop minimal question sets to elicit the user’s privacy preferences (guo2025not; tran2025understanding; zhang2025towards).
-However, as our work focuses on inferring privacy judgments on behalf of users, we leave this elicitation as a future work.
-Lastly, our work assumes a non-adversarial setting where the incoming requests are not maliciously designed.
-While this limits applicability in open environments, it is appropriate for controlled ecosystems where input sources are authenticated and vetted.
+Unfortunately, to our knowledge, there are no existing publicly-available datasets that incorporate user privacy preferences over time.
+An interesting direction for future work could elicit user privacy preferences over a period of time, rather than giving a one-time questionnaire to participants.
+Although we believe our evaluation datasets accurately reflect our problem setup, they were not originally framed as AI agents acting on the users’ behalf.
+This distinction is notable, as prior work indicates that privacy preferences can shift when information-sharing decisions are delegated to a personal AI agent (guo2025not).
+While recent studies have begun eliciting user privacy preferences regarding AI agents and chatbots (guo2025not; tran2025understanding; zhang2025autonomy; liu2025prevalence; wu2025towards; zufferey2025ai), these datasets are either (1) not publicly available, or (2) restricted to narrowly scoped scenarios.
+Consequently, this emerging field would benefit significantly from future work that appropriately releases open datasets on user privacy preferences in agentic data-sharing contexts.
+Cold-start Cases
+Our evaluation does not consider cold-start cases where the user has not made prior judgments on requests– essentially, no prior privacy preferences.
+Handling these cold-start cases is an important consideration when designing personalized agents.
+While falling back to general privacy norms is a potential solution, our results demonstrate they are often insufficient for personalized privacy decisions.
+An alternative solution could be to develop the minimal set of questions to elicit the user’s privacy preferences (guo2025privi) while offsetting user burden.
+We leave evaluation on cold-start cases as future work.
+Non-adversarial Setting
+Lastly, our work assumes a non-adversarial setting which could limit the applicability of ARIEL in open environments.
+However, ARIEL is designed as a specialized module focused on agent privacy decisions, functioning within a larger system that authenticates inputs.
+An example is Google Play Store where app developers are verified and users provide feedback on problematic apps, which prevents requests to the agent from being maliciously designed.
+Thus, defense against adversarial attacks, such as prompt injection (greshake2023not), is delegated to separate system components that focus specifically on these attacks (meng2025cellmate; tsai2025contextual).
+Additionally, ARIEL operates on limited structured input containing just five fields, inherently reducing the risk of adversarial attacks.
 8. Conclusion
 In this work, we explored personalization of LLM-based agents for making privacy judgments.
-We identified that general privacy norms are insufficient for effective personalization, and LLM-based reasoning with ICL does not guarantee judgments faithfully grounded on user’s prior privacy judgments.
-We addressed these limitations by designing ARIEL to ensure judgments are grounded in user’s prior judgments with clear and traceable reasoning that the user can reference.
-Although ARIEL is intended to accurately make privacy judgments on behalf of users, it is designed with user agency as a key component of the framework.
-Our evaluations demonstrate that ARIEL can improve appropriate judgments by reducing the F1 score error of appropriate judgments by up to 39.1% over LLM reasoning.
+We identified that general privacy norms are insufficient for effective personalization, and LLM-based reasoning with ICL does not guarantee that judgments are faithfully grounded in the user’s prior privacy judgments.
+We addressed these limitations by designing ARIEL to ensure judgments are grounded in the user’s prior judgments with clear and traceable reasoning that the user can reference.
+Our evaluations demonstrate that ARIEL can reduce the F1 score error for appropriate judgments by up to 40.6% over LLM reasoning.
 Overall, our findings suggest that jointly leveraging LLMs with logical entailment is a promising direction for designing agentic frameworks for personalized privacy judgments.
@@
 We would like to thank Adrià Gascón, Sarah Meiklejohn, Eugene Bagdasarian, Daniel Ramage, Zhichen Zhao, and Ku Kogan for their constructive discussions and support.
+The authors used generative AI-based tools to revise the text, improve flow and correct any typos, grammatical errors, and awkward phrasing.
+This work is supported by NSF award number DGE-1842487 (J.F.) and the REAL@USC-Meta center (J.F., M.A.).
 Appendix A Prompts for Evaluation